### v1

# CLARA: A Modular Framework for Unsupervised Transit Detection Using TESS Light Curves

Abstract
We present CLARA, a modular framework for unsupervised transit detection in TESS light curves, leveraging Unsupervised Random Forests (URFs) trained on synthetic datasets and guided by morphological similarity analysis.
This work addresses two core questions:
(a) How does the design of synthetic training sets affect the performance and generalization of URFs across independent TESS sectors?
(b) Do URF anomaly scores correlate with genuine astrophysical phenomena, enabling effective identification and clustering of transit-like signals?
We investigate these questions through a two-part study focused on (1) detection performance optimization, and (2) the physical interpretability of anomalies.
In Part I, we introduce three URF model variants tuned via $\alpha$-controlled scoring objectives, and evaluate their generalization across five TESS sectors. This large-scale test involved scoring 384,000 individual light curves (128,000 light curves per $\alpha$ variant), revealing stable, interpretable differences between recall-optimized, precision-optimized, and balanced models.
In Part II, our optimized clustering (DPMM Cluster 2) yields a 14.04% detection rate (16 confirmed transits among 114 candidates) from the first five TESS SPOC sectors. This reflects a substantial enrichment over baseline rates: 0.4569% for the full TESS-SPOC project candidate set (7658 candidates across 1.68 million light curves), and 0.2650% for the FFI-based SPOC sample (7658 candidates across 2.89 million light curves; Doyle et al. 2024).
All computations were performed on a personal, CPU-only desktop with an Intel® Core™ i3-8100 processor and 32 GB RAM, using parallelized scoring and classification routines across four physical cores (see Section 5). CLARA processed over 87,000 TESS SPOC light curves (Sectors 1–5) without GPU acceleration. The complete codebase is available on Github.
keywords:
software: machine learning – exoplanets – methods: data analysis – astrometry – stars: general – (stars:) binaries (including multiple): close – planet–star interactions
pubyear: 2025pagerange: CLARA: A Modular Framework for Unsupervised Transit Detection Using TESS Light CurvesD
1 Introduction
The discovery and characterization of astrophysical phenomena from large-scale time-series surveys represents one of the most challenging problems in modern astronomy. With missions like the Transiting Exoplanet Survey Satellite (TESS), Kepler, K2 and the upcoming LSST observatory generating millions of light curves, traditional supervised approaches for anomaly detection face fundamental limitations due to the rarity and diversity of interesting astrophysical events . The need for scalable, unsupervised methods that can systematically discover and interpret anomalies - such as exoplanet transits, eclipsing binaries, and novel variable stars - has become increasingly urgent as survey data volumes continue to grow exponentially.
Recent advances in unsupervised machine learning for astronomical time-series analysis have shown promising results. Nun et al. (2015) demonstrated the effectiveness of Random Forest-based approaches for stellar variability classification, while Malanchev et al. (2021) explored deep learning methods for supernova discovery in large surveys. Active learning approaches have also emerged as powerful tools for personalized anomaly detection, with Lochner & Bassett (2016) introducing frameworks that combine human intuition with machine learning efficiency. The application of Unsupervised Random Forests (URFs) to exoplanet detection was demonstrated by Crake & Martínez-Galarza (2023) (MG23 hereafter), who showed that synthetic training sets could guide anomaly detection behavior in TESS light curves, with an approach optimizing models for the sharpest anomaly score tail. From MG23:
Once the parameters that maximise this accuracy are identified, we fine-tune the specific hyperparameters near these values to increase the contrast in anomaly scores between objects.
However, these approaches have lacked systematic frameworks for controlling the discovery process and interpreting the physical nature of detected anomalies.
The challenge of anomaly detection in astrophysical time-series data is compounded by several factors: the high dimensionality of light curve features, the presence of various noise sources and systematic effects, and the need to distinguish between different classes of astrophysical phenomena Ivezić et al. (2019). Traditional clustering and dimensionality reduction techniques often fail to provide physically meaningful groupings of astronomical objects Ivezić et al. (2019), while purely data-driven approaches may discover statistically significant but astrophysically uninteresting patterns. Moreover, the analysis of time series photometry requires Lomb-Scargle periodogram techniques VanderPlas (2018). to properly characterize the temporal variability of astronomical sources that have been unevenly sampled.
We present CLARA (Controllable Learning for Anomaly Recognition in Astrophysics), a comprehensive pipeline that addresses these limitations through systematic synthetic set design and feature-to-metric mapping MG23. Our approach introduces novel methodologies for steering URF behavior through controlled variation of synthetic training parameters, enabling the targeted discovery of specific astrophysical phenomena. Furthermore, we develop morphological classification techniques using cosine similarity matching with known objects, and integrate astrometric data from Gaia DR3 Gaia Collaboration et al. (2023) to provide physical interpretation of discovered anomalies. This work aims to transform unsupervised anomaly detection from a purely statistical exercise into a physically guided discovery tool for astronomical surveys.
2 Guide for the Reader
This paper is organized into two main parts, corresponding to the two research questions outlined in the abstract.
Part I (Section 3) investigates how synthetic light curve construction influences the performance of Unsupervised Random Forest (URF) models for anomaly detection in TESS 2-minute cadence SPOC light curves. It introduces a parameterized scoring metric controlled by a variable $\alpha$ that allows users to balance detection recall and astrophysical significance. We perform a grid search over synthetic set parameters (e.g., duration, noise level, cadence) and evaluate model behavior using TOI Recall, TOI Importance AUC, and Anomaly Rate. Generalization across sectors and benchmarking against prior URF variants (URF-1 to URF-3) are also presented.
Part II (Section 4 onward) validates the astrophysical significance of high-scoring URF anomalies. We introduce a revised scoring scheme (Normalized Weighted Root Sum of Squares, or N-WRSS), and correlate anomaly scores with known stellar parameters (e.g., $T_{\mathrm{eff}}$, $\log g$, RUWE, $v_{\mathrm{tan}}$) from Gaia DR3. Morphological classification is performed using cosine similarity to labeled TOIs with SIMBAD object types (e.g., Pl, EB*, V*, PM*), allowing us to quantify the physical plausibility of detected signals.
Readers primarily interested in the machine learning framework and model tuning may focus on Part I, while those focused on astrophysical validation and interpretation may proceed directly to Part II. CLARA’s modular pipeline and evaluation methods are designed for both domains, with figures, tables, and code references included throughout to support reproducibility. For implementation details, datasets, and code repositories, please refer to the Data Availability section and the relevant Appendices.
Definitions and Terminology
We define the principal terms, abbreviations, and concepts used throughout the paper.
TESS: The Transiting Exoplanet Survey Satellite, a NASA mission that collects light curves from thousands of stars across multiple observing sectors, enabling large-scale time-domain photometry for exoplanet detection.
SPOC: Science Processing Operations Center. SPOC light curves are 2-minute cadence flux time series extracted and calibrated by NASA’s official TESS pipeline.
FFI: Full Frame Images. FFIs are 30-minute (later 10-minute and 200s) cadence image frames used to extract light curves across the entire TESS field of view. These are less targeted and more inclusive than SPOC.
TOI: TESS Object of Interest. A candidate planetary transit signal identified by the TESS pipeline and subjected to vetting.
URF: Unsupervised Random Forest. A machine learning model trained to distinguish real data from synthetic noise by learning an embedding space where outliers—defined by anomaly scores—can be identified.
Synthetic Set: A generated dataset of artificial light curves (typically noise-only or transit-injected) used to train the URF. The synthetic set acts as a pseudo-negative class for contrastive learning.
Anomaly Score: A numerical score produced by the URF that reflects how different a given light curve is from the synthetic training distribution. High scores indicate potential anomalies (e.g., transits, eclipses, instrumental artifacts).
WRSS: Weighted Root Sum Score. A statistical comparison metric used to evaluate score separability between models or datasets (e.g., between p3 and p9 URF variants).
Cosine Similarity Classification: A technique to match anomalous light curves against labeled reference examples (e.g., known eclipsing binaries or transits), using dimensionality-reduced feature vectors and cosine distance.
DTW: Dynamic Time Warping. A curve alignment method used to compare phase-folded light curves with variable durations and shifts.
Phase Folding: The process of folding a time series over a detected periodicity to align repeating patterns (e.g., planetary transits), enhancing signal-to-noise and making periodic features easier to identify.
DBSCAN: Density-Based Spatial Clustering of Applications with Noise. A clustering algorithm used to group similar anomalies and isolate dense clusters (e.g., eclipsing binaries, transit candidates).
DPMM: Dirichlet Process Mixture Model. A Bayesian non-parametric clustering technique that infers the number of clusters from the data and is used here to cluster anomaly scores or features.
Cluster -1: A reserved cluster label for outliers or noise points rejected by DBSCAN, typically enriched with rare or unique light curve morphology.
Morphological Similarity: The degree to which two light curves resemble each other in shape after normalization and alignment, used to propagate labels or identify classes.
Stratified Random Sampling: A sampling technique that divides the dataset into distinct subgroups (strata) based on a specific variable (e.g., anomaly score, sector, class label), and then samples proportionally from each group. This ensures balanced representation across the full range of a key feature or label, reducing sampling bias and preserving distributional structure.
[…]
3 Exploring how the design of synthetic sets affects unsupervised anomaly detection
Unsupervised random forest models for anomaly detection were first used in astronomical datasets in Baron & Poznanski (2017) where isolation forests were applied to SDSS galaxy data to identify the most unusual and potentially rare objects in the survey. As demonstrated ahead, the controllability of unsupervised anomaly detection hinges on our ability to systematically design synthetic training sets that guide model behavior toward discovering specific types of astrophysical phenomena. By varying key parameters such as the number of synthetic curves in the artificially generated data set, duration of curves/transits, frequency, and noise characteristics in our synthetic light curves, we can steer Unsupervised Random Forests to prioritize different anomaly signatures and effectively "tune" the discovery process for targeted scientific objectives.
3.1 Feature of a Curve
The characteristic of a single light curve for input to our URF (Unsupervised Random Forest) model experiments follows established methods from our foundational paper MG23 where the first n=3000 flux points are stacked on top of nls = 1000 Lomb-Scargle periodogram, VanderPlas (2018) frequency points calculated using the LombScargle function of the astropy.timeseries module. The feature is thus an array with 4000 values. This function calculates the power of a time period where the power describes the extent of the periodicity of the selected time value, given by:
$P_{LS}(\omega)=\frac{1}{2}\left[\frac{\left(\sum_{j}(x_{j}-\bar{x})\cos(\omega(t_{j}-\tau))\right)^{2}}{\sum_{j}\cos^{2}(\omega(t_{j}-\tau))}\right.\\ \left.+\frac{\left(\sum_{j}(x_{j}-\bar{x})\sin(\omega(t_{j}-\tau))\right)^{2}}{\sum_{j}\sin^{2}(\omega(t_{j}-\tau))}\right]$ (1)
where:
$P_{LS}(\omega)$ $=\text{Lomb-Scargle power at angular frequency }\omega$
$x_{j}$ $=\text{observed flux values at times }t_{j}$
$\bar{x}$ $=\text{mean of the observed flux values}$
$\tau$ $=\text{time offset defined by:}$
$\tan(2\omega\tau)$ $=\frac{\sum_{j}\sin(2\omega t_{j})}{\sum_{j}\cos(2\omega t_{j})}$
An interesting property of the Lomb-Scargle periodogram is that it can, in principle, fail to detect strictly periodic signals if the sampling intervals are strictly irrational. However, this is more of a mathematical curiosity than a practical concern. In real-world observations, sampling intervals are always rational (due to finite precision) and periodic signals, even if originally irrational in frequency, are effectively rationalized by floating-point limitations in computation. Additionally, computing a Nyquist frequency becomes ill-defined under irrational sampling intervals, further complicating standard aliasing interpretations - but again, this rarely arises in empirical datasets.
Figure 1: The first 3000 flux points of a light curve (top plot) and the Lomb-Scargle periodogram of the same curve (bottom plot)
3.2 URF Hyperparamter Randomized Search via Real + Synthetic Curve Stacking
For the real set (real features class), we took features of 1500 real curves from sector 2. A RandomizedSearchCV is performed over a predefined hyperparameter space for the RandomForestClassifier. The search is:
  • Randomized (n_iter=30)
  • Parallel (n_jobs=-1)
  • Reproducible (random_state=42)
The hyper-parameter search space is mentioned in Appendix A.2
For our initial tests, we took the following synthetic set (artificially generated features class) design types for the URF.
a. URF1. We generated the synthetic set by uniformly sampling feature values within the range of the real dataset. This establishes a baseline benchmark method.
X_fake = np.random.uniform(np.min(X_real), np.max(X_real), size=X_real.shape)
b. URF2. We used fixed hyper parameters mentioned in MG23. The following hyper parameters were used:
  • n_estimators = 100
  • max_features = "sqrt"
  • max_depth = 700
  • min_samples_split = 4
  • min_samples_leaf = 2
  • bootstrap = False
  • warm_start = True
c. URF3. We used a TOI based synthetic class where we obtained feature vectors from Tess Objects of Interest (TOIs) from Sector 1 of SPOC data, and set those features as the synthetic set for hyper-parameter search.
d. URF4. The synthetic set of features were computed from a set of synthetic light curves generated using transit models (box-shaped model and Mandel-Agol) injected into realistic noise baselines. The computational implementation of this was performed using the batman pypi package.
The four URF variants were designed to systematically explore different approaches to synthetic set construction for transit discovery. URF1 serves as a baseline control, using uniform random sampling within the feature space bounds to establish minimum performance expectations. URF2 tests the generalizability of the hyper parameters established in MG23, evaluating whether their optimization for stellar variability classification translates effectively to transit detection in TESS data. URF3 explores a realistic "ground truth" approach by using feature vectors from confirmed TOIs as the synthetic class, providing insight into the performance ceiling when the synthetic set closely matches actual transit signatures. Finally, URF4 introduces our controllable framework, where synthetic light curves are generated using physically motivated transit models with systematically variable parameters (ratio of real curves - 1500 - to synthetic curves - 300), duration, cadence, and noise characteristics), enabling targeted optimization for specific discovery objectives and forming the foundation for our systematic parameter exploration. For URF4, we fixed the real set to 1500 TESS SPOC curves across all experiments; investigating how the size and composition of this real subset affects model performance is a valuable direction for future work.
3.3 Anomaly Scoring
Following Baron & Poznanski (2017) and MG23, our URF implementation uses terminal node population as the anomaly scoring heuristic. Each tree in the forest is trained to distinguish real light curves from synthetic ones, with anomaly scores computed by tracking the population composition of terminal nodes. For each real light curve, we calculate a similarity score S as the average fraction of real data points sharing the same terminal node across all trees in the forest. The final anomaly score is defined as 1 - S, where a score of 1 indicates maximum anomaly (the object consistently lands alone in terminal nodes) and 0 indicates minimal anomaly (the object is always grouped with the majority of real data). This population-based approach provides a natural measure of how isolated each light curve appears relative to the broader dataset distribution.
Table 1: Performance comparison of URF variants on TOI detection. TP = True Positives, FP = False Positives, TN = True Negatives, with precision and recall calculated for the 175 total TOIs in the test set for URFs 1,2 and 4 and 195 total TOIs in the test set for URF 3
URF Variant TP FP TP+FP TN Precision (%) Recall (%)
URF-1 22 2159 2181 0 1.01 12.57
URF-2 45 3990 4035 0 1.12 25.71
URF-3 14 300 314 15499 4.46 7.18
URF-4 64 4982 5046 0 1.27 36.57
3.4 Initial Results
Table 1 summarizes the performance of our four URF variants on TOI detection across the test dataset. URF-1, URF-2, and URF-4 demonstrate similar scoring behavior, assigning zero scores to most light curves and treating only non-zero scores as anomalies. URF-1 and URF-2 achieve baseline recalls of 12.6% and 25.7% respectively with 1̃% precision, while URF-4 shows improved performance at 36.6% recall and 1.3% precision. URF-3 exhibits fundamentally different behavior, assigning non-zero scores to nearly all light curves, requiring threshold-based evaluation (scores > 0.22, chosen arbitrarily based on figure 3) to define meaningful anomaly regions. When evaluated under these conditions, URF-3 achieves 7.2% recall with 4.5% precision, though this performance is best assessed on independent sectors to avoid overfitting bias from using TOI features as synthetic training data. The distinct scoring profiles highlight URF-4’s advantage in providing both strong recall performance and systematic controllability without the threshold sensitivity issues observed in URF-3
Figure 2: URF3 TOI (top) ranging from 0.0 to 0.65 with a normal shaped distribution ($\mu\approx$ 0.45) and All Curves (bottom) Anomaly Score with a normal shaped distribution ($mu\approx 0.1$) for Sector 1 Test Set showing distinct background separation between TOI anomaly scores and non TOI anomaly scores, interpreted as a result of URF overfitting to sector 1 TOI features as proved by same distribution plotted for sector 2 test set as in Figure 3
Figure 3: URF3 TOI (top) normal shaped distribution ($\mu\approx 0.1$) and All Curves (bottom) Anomaly Score normal shaped distribution ($\mu\approx 0.1$) for Sector 2 Test Set, showing TOI distribution (top) with near perfect distribution value and shape overlap over the entire anomaly score distribution (bottom)
URF-3 showed significant overfitting to TOI curve features of sector 1 which resulted in the distinct separation of TOIs from normal curve background when URF3 scored on sector 1. This was proved by cross sector evaluation of URF3 on SPOC sector 2 curve set. Refer to appendix A.1
3.5 Controlled Parameter Space Exploration
We selected URF-4 for comprehensive parameter optimization based on its superior recall performance (36.6%) during our initial benchmarking evaluation (URF 1-4) and its inherent controllability through physically motivated synthetic generation. URF-4’s framework enables systematic exploration of four key synthetic parameters (and more not explored ones):
  1. 1.
    ncurves - the number of synthetic light curves in the artificially generated training set
  2. 2.
    duration - temporal span of each synthetic curve in days
  3. 3.
    cadence minutes - observational temporal resolution
  4. 4.
    noise level - injected photometric noise in parts-per-million (ppm)
The following parameter grid was used for this purpose:
    param_grid = {
    ’n_curves’: [100, 200, 300],
    ’lc_length_days’: [13.5, 27.0],
    ’cadence_minutes’: [2.0, 10.0],
    ’noise_level’: [50e-6, 100e-6, 200e-6],
Our analysis employs a systematic parameter sweep methodology, generating synthetic training sets for a URF4 sub-variant model by varying individual parameters while holding others constant at baseline values. Each parameter combination from the cartesian product space is used to train (3.2) independent URF models against a fixed real light curve dataset, resulting in $3\times 2\times 2\times 3=36$ sub-variant models.
3.6 Evaluation Methodology
To ensure robust and unbiased assessment of our parameter sweep results, we implement a controlled evaluation framework with two key components: representative test set construction and consistent subset sampling for fair model comparison.
3.6.1 Representative Test Set Construction
We construct our test set using stratified random sampling to maintain the same TOI-to-total light curve ratio as observed in the full sector data. This approach ensures that performance metrics remain meaningful and are not artificially inflated or deflated by over- or under-representation of confirmed TOIs in the evaluation set. The stratified sampling preserves the natural class distribution, enabling fair comparison of model performance against realistic detection scenarios.
3.6.2 Controlled Subset for Model Comparison
To enable computationally efficient and methodologically sound comparison across all 36 URF sub-variant models, we employ a fixed subset approach for both training and evaluation. A randomly selected stratified subset of 4,000 light curves is drawn from the full SPOC 2-minute cadence sector 2 dataset and used consistently across all parameter combinations. This controlled framework offers several advantages:
  • Computational efficiency: Reduces training and evaluation time across the extensive parameter grid
  • Controlled comparison: Eliminates variation due to different light curve content across experimental runs
  • Isolated parameter effects: Ensures performance differences are attributable solely to synthetic set design choices rather than random sampling variation
  • Trend Analysis Step: The performance of the 36 models on this small subset helps us analyze trends in their behavior in relation to the parameters of synthetic set generation and quantify those trends using the performance metrics in 3.6.4 and then test the generalization of those trends on independent sectors using cross sector evaluation of the models with more (and better, in terms of representation of quantification of model performance) performance metrics described in Table 3
3.6.3 TOI Importance Weighting
To prioritize TOIs by their detectability and astrophysical significance, we define an importance score that favors targets with stronger observational signatures:
$\text{Importance Score}=\text{SNR}\times\text{Depth (ppm)}\times\text{Duration (days)}$ (2)
where SNR represents the signal-to-noise ratio of the planet detection, Depth (ppm) is the transit depth in parts per million, and Duration (days) is the transit duration in days. This heuristic emphasizes TOIs that are deeper, longer-duration, and higher signal-to-noise, making them both more detectable and more reliable for anomaly detection evaluation.
For fair comparison across sectors, TOI importance scores are normalized to the range $[0,1]$:
$\text{Normalized Importance}=\frac{\text{Importance Score}-\text{Score}_{\text{min}}}{\text{Score}_{\text{max}}-\text{Score}_{\text{min}}}$ (3)
This normalization enables sector-independent comparison, with higher normalized scores indicating greater relative astrophysical importance within the evaluation set. The TOI Importance AUC metric then rewards models that assign high anomaly scores to the most observationally significant targets, providing a more nuanced assessment than simple recall alone.
3.6.4 Performance Metrics
We evaluate URF model performance using three complementary metrics designed to capture both detection efficiency and astrophysical relevance:
  • Anomaly Rate: Fraction of the test set flagged as anomalous, providing insight into model selectivity
  • TOI Recall: Fraction of known TOIs successfully recovered, measuring detection completeness
  • TOI Importance AUC: Area under the curve when TOIs are ranked by anomaly score and weighted by astrophysical importance as described in section 3.6.3
Threshold-Based Recall Analysis: To assess how effectively each URF-4 sub-variant ranks TOIs among detected anomalies, we perform systematic threshold analysis across varying anomaly score cutoffs. For each threshold corresponding to the top $n\%$ of anomaly scores (where $n\in\{5,10,15,\ldots,100\}$), we calculate:
[…]
This analysis reveals whether TOIs are preferentially concentrated at higher anomaly scores, indicating effective ranking behavior, and quantifies how recall performance varies as detection thresholds are adjusted.
It is worth noting that TOI Importance AUC (and other importance metrics derived further on) and TOI Recall serve as astronomy-contextualized analogs of traditional precision-recall metrics in machine learning performance analysis. While standard recall measures the fraction of positive class instances correctly identified, TOI Recall specifically quantifies the fraction of confirmed/nominated transit candidates recovered by the anomaly detection framework. Similarly, TOI Importance AUC extends beyond simple area-under-curve calculations by weighting detections according to their observational significance, providing a domain-specific assessment of model ranking quality that prioritizes the most scientifically valuable targets.
3.7 Correlating Synthetic Parameters to Performance Metrics
3.7.1 Results (1)
Here, we present the systematic trends we observed for these 36 urf-4 sub-variant models. Figures 4, and 5 present the correlation between individual synthetic parameters and the performance metrics TOI Recall, Anomaly Rate and TOI Importance AUC (A.4).
Figure 4: ncurves, duration_hr (days A.3), noise (ppm) and cadence (minutes) clockwise from top left against TOI Recall (blue) and Anomaly Rate (red)
Figure 5: Synthetic features (ncurves, duration_hr(in days A.3), noise (ppm), cadence (minutes) - clockwise from top left) v.s TOI Importance AUC
To quantify the relative influence of synthetic training parameters on URF-4 performance, we employed Random Forest regression to model the relationship between our four key synthetic parameters (n_transits, duration_hr A.3, cadence_min, noise_ppm) and three primary evaluation metrics (TOI Recall, Anomaly Rate and TOI Importance AUC). This approach provides interpretable feature importance scores that reveal which synthetic design choices most strongly influence model behavior.
The Random Forest regression analysis, used to quantify how synthetic training parameters influenced URF-4 performance (Figure 6, 7), directly informed the finer-grained experiments presented in Section 4.1. In particular, noise (ppm) emerged as the most influential parameter affecting TOI Importance AUC and transit duration the most influential for anomaly rate and TOI Recall behavior. Motivated by this, we fixed other parameters and systematically varied noise and duration to study their specific impact on multiple performance metrics using a revised, multi-component scoring formulation 4.1 for $\alpha$ model selection for experiments in Section 4. This highlights how feature importance insights can guide targeted parameter space explorations and model selection strategies.
Figure 6: Feature importance comparison for URF-4 synthetic training parameters. Left panel shows importance scores for TOI Recall prediction, right panel for Anomaly Rate prediction. Transit duration (days A.3) emerges as the dominant factor for both metrics, followed by transit multiplicity. Random Forest regressors trained on 36-variant parameter sweep results with 42-seed random state for reproducibility.
Figure 7: Feature importance comparison for URF-4 synthetic training parameters. This plots TOI Importance AUC feature importance against the features.
Figure 8: Performance of top 10 URF-4 variants across anomaly threshold percentiles. Left panel shows TOI Recall AUC curves for the highest-performing parameter configurations. Right panel presents TOI Importance AUC trajectories.
Figure 8 presents the performance trajectories of the top 10 URF-4 variants ranked by TOI Recall AUC across varying anomaly threshold percentiles. It shows the top URF-4 optimal models achieving  40% recall in the top 25% of anomaly scores. However, while individual feature importance rankings guide parameter prioritization for focused optimization (4), the combined metric analysis 3.8 becomes essential for interpreting current results since different parameters optimize different performance aspects. The divergent trajectories between TOI Recall AUC and TOI Importance AUC in Figure 8 highlight the need for weighted combination scoring to balance detection completeness against astrophysical relevance in variant selection.
3.8 Combined Metric Analysis
To optimize synthetic light curve parameters for specific scientific objectives, we introduce a combined performance metric that balances detection completeness against astrophysical relevance:
$\text{S}_{\text{combined}}=\alpha\cdot\text{Recall AUC}_{\text{TOI}}+(1-\alpha)\cdot\text{Importance AUC}_{\text{TOI}}$ (5)
The weighting parameter $\alpha$ enables systematic control over the scientific trade-off between casting a wider net (high $\alpha$, emphasizing comprehensive TOI recall) versus casting a more accurate net (low $\alpha$, prioritizing astrophysically significant (higher values of importance metrics) targets). This parameterization allows researchers to tailor anomaly detection behavior to their specific discovery goals: survey completeness studies favor high $\alpha$ values to maximize detection rates, while targeted follow-up campaigns benefit from low $\alpha$ values that concentrate observational resources on the most promising candidates.
For a given $\alpha$, the combined score $\text{S}_{\text{combined}}$ is calculated for each urf-4 variant model and then ranked to select the top model.
This formulation satisfies the following properties:
  1. 1.
    Convexity: Since $\alpha\in[0,1]$ $S_{\text{combined}}$ lies within the convex hull of $R$ and $I$. This guarantees that the combined score is always bounded between the two metrics and cannot extrapolate beyond them i.e. $S_{\text{combined}}\in[\min(R,I),\max(R,I)]$.
  2. 2.
    Continuity: The score varies smoothly as $\alpha$ changes, enabling gradient-free optimization (e.g., grid search) over $\alpha$ values.
  3. 3.
    Trade-off Interpretation:
    • $\alpha\to 1$ emphasizes Recall AUC, optimizing for high completeness (i.e., casting a wide net to detect as many TOIs as possible).
    • $\alpha\to 0$ emphasizes Importance AUC, favoring models that score astrophysically significant TOIs highly, even at the cost of total recall. A more comprehensive counterpart of this combined scoring metric has been presented in Equation 4.1.
In the second part of this research 4, we use the intersection of high alpha and low alpha models and use an improved anomaly scoring method - Normalized Weighted Room Sum Score (N-WRSS) - for quantifying anomaly candidates.
3.8.1 Generalization Evaluation of Balanced URF Variant ($\alpha$=0.5)
To validate the stability and generalization of our optimal URF-4 configuration for $\alpha$=0.5, we evaluate the balanced model across multiple random subsets of unseen data compiled with stratified sampling from the same sector (sector 2). In this context, by "generalization" we refer to how consistent a model’s performance is across 10 randomly sampled, stratified subsets of the same sector, that is, its stability over similar data distributions. Specifically, we compute the mean and deviation of performance values relative to predictions made from the trend-prediction test set defined in Section 3.6.2.
Selected Model ($\alpha=0.5$): urf4_model_variant_n200_d13_c2_n50ppm
Target Performance Metrics (trend analysis results):
  • TOI Recall AUC $\approx$ 0.561
  • TOI Importance AUC $\approx$ 0.120
Evaluation Protocol: Our generalization assessment employs a rigorous random sampling approach:
  1. 1.
    Random Sampling: Generate 10 disjoint/overlapping subsets of 4,000 light curves each from the full sector 2 SPOC dataset.
  2. 2.
    Anomaly Scoring: Apply the selected URF model to compute anomaly scores for each subset.
  3. 3.
    Subset Analysis: For each subset, compute TOI Recall AUC and TOI Importance AUC using known TOI classifications.
  4. 4.
    Statistical Aggregation: Calculate mean and deviation of mean from predicted value across all subsets to assess stability.
Pass Criteria: A model demonstrates acceptable generalization if:
  • Both metrics exhibit low variance (standard deviation < 10% of mean)
  • Mean performance remains within ±0.03–0.05 of target values
Results: Evaluation across 10 random subsets ($\approx$40,000 total light curves) yields Table 2
Metric Value Target Deviation
Mean TOI Recall AUC 0.547 0.561 -0.014
Mean TOI Importance AUC 0.143 0.120 +0.023
Mean Binary TOI Recall 0.321 - -
Mean Anomaly Rate 0.279 - -
Table 2: Generalization performance of balanced URF-4 variant ($\alpha$ = 0.5)
Metric What It Measures Why It Matters
TOI Recall AUC Distribution quality of TOIs toward top anomaly scores Reflects overall prioritization quality across the full score spectrum
TOI Importance AUC Preservation of high-value TOIs across score thresholds Indicates whether scientifically valuable objects are ranked highly
Mean Importance (Top 20%) Average importance of TOIs within top 20% of scores Shows strength of early-ranked candidates - critical for targeted follow-up
Binary TOI Recall Fraction of TOIs flagged as anomalies (score > 0) Direct measure of detection breadth for candidate mining
Anomaly Rate Fraction of all light curves flagged as anomalous Reflects follow-up cost - high rates imply more manual vetting
Percentile Range (Top N%) Anomaly score percentile range containing top N% of TOIs by Importance Score Measures how early high-value objects appear - useful for threshold setting
Table 3: Six-metric evaluation framework for $\alpha$-dependent behavior analysis
Interpretation: The evaluation demonstrates strong model generalization: TOI Recall AUC deviates only 0.014 below target, indicating consistent TOI prioritization across the anomaly score distribution. TOI Importance AUC exceeds the target by 0.023, suggesting the model performs even better than expected at detecting astrophysically significant targets. The binary recall of 32.1% and anomaly rate of 27.9% indicate closely related TOI Recall and Anomaly Rate as predicted by Figure 4. The value of these metrics will be improved by predicting over a finer mesh of synthetic design features and using a more comprehensive model scoring metric as presented in Section 4.1.
Verdict: This model satisfies our generalization criteria (that it will perform consistently on similar data), with both metrics falling within the acceptable ±0.03-0.05 deviation range. The empirical results confirm that the $\alpha$ = 0.5 design successfully balances recall and importance while maintaining stable performance across similar (in terms of ratio of TOIs to total curves) randomly stratified sampled data subsets from sector 2.
3.8.2 Hypothesis Validation: Alpha Variants
To validate the hypothesis that synthetic set design can be used to steer model behavior toward different scientific objectives, two additional URF-4 subvariants are evaluated over 10 sets of 4,000 light curves each - totaling 40,000 light curves per model (the same subsets used to evaluate the balanced model with $\alpha$ = 0.5) - with configurations selected based on $\alpha$-value ranking:
  • Low-$\alpha$ model ($\alpha$ = 0.3): Emphasizes astrophysical importance
  • High-$\alpha$ model ($\alpha$ = 0.9): Prioritizes detection completeness
Goal: Assess whether changing $\alpha$ systematically alters model behavior across multiple evaluation dimensions - not just recall versus importance, but also prioritization sharpness, anomaly rate, and TOI distribution in score space. Model selection by this $\alpha$ metric is subject to a light supervisory step as referred to in Appendix A.5.
Evaluation Framework: Six-Metric Assessment
URF-4 $\alpha$ variants are evaluated across six complementary metrics, each capturing different aspects of anomaly detection performance and scientific utility as in Table 3
3.8.3 Results: Alpha Variant Performance Comparison
Evaluation across 10 consistent test subsets (40,000 total light curves per model) reveals clear and reproducible performance differences between $\alpha$ variants, as summarized in Table 4.
Metric $\alpha$ = 0.3 $\alpha$ = 0.5 $\alpha$ = 0.9 Best Performer
TOI Recall AUC 0.521 0.547 0.557 $\alpha$ = 0.9
TOI Importance AUC 0.141 0.143 0.139 $\alpha$ = 0.5
Mean TOI Importance (Top 20%) 0.393 0.422 0.359 $\alpha$ = 0.5
Binary TOI Recall 0.493 0.321 0.682 $\alpha$ = 0.9
Anomaly Rate 0.447 0.279 0.558 $\alpha$ = 0.5
Top 10% Importance TOIs – Score Percentile Range 1.6–9.7% 4.9–10.6% 2.0–10.2% $\alpha$ = 0.3
Top 20% Importance TOIs – Score Percentile Range 1.3–28.4% 3.1–24.4% 1.4–23.8% $\alpha$ = 0.3
Table 4: Performance comparison of URF-4 $\alpha$ variants across six evaluation metrics
Metric S1 Best S2 Best S3 Best S4 Best S5 Best Consistency
TOI Recall AUC $\alpha$=0.5 $\alpha$=0.9 $\alpha$=0.9 $\alpha$=0.9 $\alpha$=0.5 60% $\alpha$=0.9
TOI Importance AUC $\alpha$=0.9 $\alpha$=0.5 $\alpha$=0.5 $\alpha$=0.3 $\alpha$=0.3 Mixed
TOI Importance Top 20% $\alpha$=0.5 $\alpha$=0.5 $\alpha$=0.5 $\alpha$=0.5 $\alpha$=0.5 100% $\alpha$=0.5
Binary TOI Recall $\alpha$=0.9 $\alpha$=0.9 $\alpha$=0.9 $\alpha$=0.9 $\alpha$=0.9 100% $\alpha$=0.9
Anomaly Rate (lowest) $\alpha$=0.5 $\alpha$=0.5 $\alpha$=0.5 $\alpha$=0.5 $\alpha$=0.5 100% $\alpha$=0.5
Top 10P Importance Range $\alpha$=0.3 $\alpha$=0.3 $\alpha$=0.3 $\alpha$=0.5 $\alpha$=0.5 60% $\alpha$=0.3
Top 20P Importance Range $\alpha$=0.3 $\alpha$=0.9 $\alpha$=0.5 $\alpha$=0.5 $\alpha$=0.9 Mixed
Table 5: Cross-sector performance of $\alpha$ variants across TESS Sectors 1-5
3.8.4 Intepretation of Results of Table 4
Behavioral Analysis: Systematic $\alpha$-Dependent Performance - The three $\alpha$-variant models demonstrate consistent and distinct behavioral patterns that align with their design objectives:
$\alpha$ = 0.9 (High Recall Priority): Emphasizes detection completeness, achieving the highest TOI Recall AUC (0.557) and Binary TOI Recall (68.2%). However, this comes at the cost of an elevated anomaly rate (55.8%), indicating increased false positive burden for follow-up observations.
$\alpha$ = 0.3 (High Importance Priority): Optimizes early ranking quality, concentrating the most scientifically valuable TOIs into narrow score percentile ranges (top 10% importance TOIs span only 1.6–9.7% of score space). This model achieves effective importance-based prioritization while maintaining moderate recall performance.
$\alpha$ = 0.5 (Balanced Priority): Provides an optimal trade-off, achieving the lowest anomaly rate (27.9%) while maintaining strong performance across importance metrics. The Mean TOI Importance in the top 20% reaches 0.422, indicating effective early detection of high-value targets with minimal false positive contamination.
These results validate the central hypothesis: synthetic set design successfully steers model behavior toward different scientific objectives. Each $\alpha$ variant exhibits predictable performance characteristics that remain stable across randomized subsets, demonstrating that:
  1. 1.
    $\alpha$-parameter control works as designed: Higher $\alpha$ values consistently increase recall-based metrics while lower $\alpha$ values enhance importance-based prioritization
  2. 2.
    Behavioral differences are robust: Performance patterns persist across all 10 evaluation subsets
  3. 3.
    Trade-offs are quantifiable: The six-metric framework captures the nuanced relationships between completeness, precision, and scientific value
3.8.5 Cross-Sector Generalization of Alpha Variants
Having established that $\alpha$ variants exhibit distinct and stable behavior within Sector 2, we next evaluate cross-sector generalization to test whether these behavioral patterns persist across different TESS observational periods. This analysis addresses a critical question for practical deployment: do models selected and characterized on one sector maintain their performance characteristics when applied to entirely different stellar populations and observing conditions?
Evaluation Strategy: Each $\alpha$ variant (0.3, 0.5, 0.9) is evaluated on the same six-metrics tracked across 5 TESS SPOC sectors S1, S2, S3, S4 and S5 to establish:
  1. 1.
    Performance stability: Whether relative $\alpha$-dependent behaviors remain consistent across sectors
  2. 2.
    Best performer identification: Which $\alpha$ value optimizes each performance metric for different stellar populations
For each sector, 10 random stratified subsets for Sectors 1 and 2, and 4 subsets for Sectors 3, 4, and 5, sample sets of 4000 light curves each were evaluated using the three $\alpha$-variant models. Sector-level means of the six performance metrics were then computed for comparative analysis. In total, each model variant scored exactly 128,000 light curves across the five sectors, resulting in an aggregate of 384,000 curves scored by an urf model across all $\alpha$ variants for this cross-sector generalization test.
The means of each of the six metrics have been presented in Figure 9. The best performing alpha variant across the metrics have been presented in Table 5.
3.8.6 Results: Cross-Sector Generalization of Alpha Variants
Cross-sector evaluation across TESS Sectors 1-5 demonstrates robust generalization of $\alpha$-dependent behavioral patterns, validating the transferability of synthetic training across diverse stellar populations.
Performance Consistency: Each $\alpha$ variant maintains its designed priorities across all sectors:
  • $\alpha$ = 0.9: Achieves best Binary TOI Recall in all 5 sectors (100% consistency) with Mean Recall ranging from 0.509 to 0.762 and best TOI Recall AUC in 3/5 sectors (60% consistency).
  • $\alpha$ = 0.3: Wins Top 10% Importance Range in 3/5 sectors (60% consistency), consistently concentrating high-value targets in narrow score ranges
  • $\alpha$ = 0.5: Achieves best TOI Importance Top 20% and lowest Anomaly Rate in all 5 sectors (100% consistency each) with TOI Importance Top 20P ranging from 0.299 to 0.486 and Anomaly Rate ranging from 0.269 to 0.311 across the five sectors.
Operational Implications: The predictable cross-sector behavior enables confident deployment of $\alpha$ variants to new TESS sectors without retraining, supporting mission-critical survey strategy decisions based on scientific priorities (set using a custom combined scoring system using performance metrics that represent that specific goals).
Figure 9: Alpha variants cross sector test performance metrics values
3.9 Benchmarking URF Models
To evaluate the impact of controllable synthetic design, we benchmark the performance of our URF-4 $\alpha$-variants against previously established baselines: URF-1 (uniform noise), URF-2 (fixed hyperparameters from MG23), and URF-3 (TOI-based synthetic set). This benchmarking provides a comprehensive view of the performance landscape for URF-based anomaly detection and situates CLARA’s synthetic design approach within it.
Limitations of Prior URF Variants:
  • URF-1: Overly conservative, achieving only 11% TOI recall. This model prioritizes purity but fails to identify most real transits.
  • URF-2: Based on hyperparameter settings from prior work. It demonstrates reasonable recall but lacks both flexibility and interpretability.
  • URF-3: Overly aggressive, labeling nearly 96% of light curves as anomalous. While it captures nearly all TOIs, it does so at the cost of overwhelming noise.
Advantage of URF-4 Synthetic Design: URF-4 circumvents these trade-offs by using physically motivated synthetic injections to enable $\alpha$-tunable control. Each $\alpha$ value corresponds to a scientific objective:
  • Alpha 0.3: Prioritizes importance-weighted metrics, ideal for high-confidence discovery pipelines.
  • Alpha 0.5: Provides an optimal balance—yielding 33% recall with only 27% anomaly rate..
  • Alpha 0.9: Maximizes completeness with high recall, suitable when follow-up resources are abundant.
Figure 10 shows a radar plot for Sector 1 comparing URF variants across normalized metrics including TOI recall, anomaly rate, importance-weighted metrics, and variability in top-importance scores. Ideal models minimize top_10p_by_importance_percentile_range_upper, strike a trade-off between toi_recall and anomaly_rate, and maximize mean_importance_top20p.
Figure 10: Normalized metric comparison across URF variants for Sector 1. URF-4 $\alpha$ models outperform fixed baselines and allow strategic tuning of performance based on discovery goals.
Key Finding: Fixed hyperparameter models (URF-1, URF-2, URF-3) lock us into rigid, suboptimal behaviors. In contrast, our synthetic design enables targeted optimization for different scientific objectives - whether maximizing recall, minimizing false positives, or improving the astrophysical quality of discoveries. URF-4 $\alpha$-variants occupy the “Goldilocks zone” of anomaly detection: not too conservative, not too aggressive, but controllably optimized and generalizable over multiple sectors.
3.10 Conclusion (Part I)
The comprehensive evaluation of URF-4 $\alpha$ variants establishes the successful validation of our central hypothesis: synthetic set design enables systematic steering of anomaly detection models toward different scientific objectives. Cross-sector analysis across TESS Sectors 1-5 demonstrates robust behavioral consistency, with each $\alpha$ variant maintaining its designed priorities across diverse stellar populations and observing conditions.
Key findings include:
  • High-$\alpha$ models ($\alpha$ = 0.9) consistently maximize detection completeness, achieving best Binary TOI Recall across all sectors
  • Low-$\alpha$ models ($\alpha$ = 0.3) excel at importance-based prioritization, concentrating high-value targets within narrow anomaly score ranges
  • Balanced models ($\alpha$ = 0.5) provide optimal precision-recall trade-offs while maintaining the lowest false positive rates.
The predictable cross-sector generalization validates the transferability of synthetic training approaches and enables confident deployment to new TESS sectors without retraining.
Having established optimal URF architectures and validated behavioral steering mechanisms, we now proceed to comprehensive parameter space optimization to identify the single best-performing model configurations for $\alpha$=0.3 and $\alpha$=0.9. Through systematic evaluation across the following parameter grid:
  • Noise levels (ppm): [25, 50, 75, 100, 125, 150, 175, 200, 225, 250, 275, 300, 400, 600, 800]
  • Duration (days A.3): [0.73, 1.65, 3.0, 3.58, 5.1, 6.0, 7.0, 8.3, 10.0, 12.0, 13.0, 15.67, 17.0, 20.0, 21.0, 21.5, 22.0, 22.5, 23.0, 23.25, 23.5, 23.75, 24.0, 24.25, 24.5, 24.75, 25.0, 25.25, 25.5, 26.0, 27.0]
  • Fixed parameters: n_curves = 300, cadence_minutes = 2
The next part of this paper validates the astrophysical significance of the optimized URF model through correlation analysis between anomaly scores and known stellar/planetary properties. Using improved Normalized Weighted Root Sum of Squares (N-WRSS) scoring and t-SNE clustering analysis, we demonstrate that high anomaly scores correspond to genuine astrophysical phenomena rather than instrumental artifacts. Morphological classification pipelines filter light curves based on SIMBAD catalog similarity, providing quantitative validation of URF detection capabilities.
4 Astrophysical Relevance and Morphological Clustering of URF Anomalies For Transit Detection
In this section, we shift focus from model selection and parameter optimization to the astrophysical validation and interpretability of anomalies detected by the CLARA framework. Having established that synthetic set design can systematically steer URF behavior, we now seek to answer a fundamental question: Do the anomalies identified by our optimized models correspond to genuine astrophysical phenomena, or are they dominated by noise and instrumental artifacts?
To address this first, we perform a feature-to-metric mapping over a finer mesh of synthetic set feature values as described in section 3.10, systematically varying synthetic light curve parameters (such as noise, and duration, as they were the most important feature according to the Random Forest Regressor) to pick 0.3 and 0.9 $\alpha$ variant urf-4 models as described in the previous section.
Second, we rigorously validate the astrophysical significance of detected anomalies. By correlating anomaly scores with known stellar and planetary properties, and employing techniques such as Normalized Weighted Root Sum of Squares (N-WRSS) scoring and t-SNE clustering (van der Maaten & Hinton, 2008), we map the distribution of high-scoring anomalies in both feature and astrophysical parameter space. Furthermore, we leverage external catalogs such as SIMBAD (Wenger et al., 2000) to perform morphological classification, quantitatively demonstrating that our models are sensitive to real transit-like events and astrophysically meaningful outliers.
Together, these analyses establish a robust framework for interpreting and validating unsupervised anomaly detection in large-scale time-series surveys, providing both scientific insight and a blueprint for extending CLARA to new domains and datasets.
4.1 A Finer Mapping of Features to Performance Metrics
Figure 11: TOI Recall (blue) vs Anomaly Rate (red) by Noise (ppm) with values [25, 50, 75, 100, 125, 150, 175, 200, 225, 250, 275, 300, 400, 600, 800]
Figure 12: Mean Importance Top 20P and Anomaly Score Range Upper 10P vs Noise (ppm)
These two parameters — noise (ppm) and transit duration — were selected for finer-grained grid variation based on feature importance analysis from earlier URF-4 performance modeling (Figures 67). Random Forest regressors revealed that noise most strongly influenced TOI Importance AUC, while duration was the dominant driver of anomaly rate and TOI Recall behavior. This empirical insight directly motivated our decision to fix all other variables and systematically vary only these two dimensions, enabling targeted exploration of their isolated effects on performance under the multi-metric $\alpha$-model selection criteria in Section 4.
The grid was intentionally densified in the ranges of duration $\sim$23–25.5 days and noise $\sim$150–300 ppm, where prior experiments indicated sharp performance transitions, enabling finer resolution of model behavior in these critical regions.
With fixed values of ncurves=300, duration(days A.3)=13 and cadence_minutes=2, we systematically varied the noise parameter in the range [25, 800]. To rank models based on varied values of noise (ppm), we used the following revised combined scoring metric:
$\text{Combined Score}=\alpha\cdot R+(1-\alpha)\cdot(1-I)-\beta\cdot A$ (6)
where a high value of $\beta$ penalizes anomaly rate $A$ more. The $\beta$ term was introduced due to significant differences between anomaly rate and recall at several values for noise as demonstrated in Figure 11. (1I)(1-I) to minimize $I$, which represents the upper limit of Top 10P Importance Range (range of anomaly score that contains the top 10 percent TOIs by importance score).
Noise (ppm) values 50e-6 and 225e-6 turned out most dominant in terms of importance. The 10Pimportance Upper Limit values for the most dominant noise values were 1.70 and 34.22 respectively. The TOI Recall values were 0.26 and 0.63 respectively. The Anomaly Rates were 0.29 and 0.58 respectively. We now set noise (ppm) at 50 and 225 and move on to varying values of duration, with more values in the 13 days to 27 days range. The following values were tested for duration (days A.3):
0.73, 1.65, 3.0, 3.58, 5.1, 6.0, 7.0, 8.3, 10.0, 12.0, 13.0, 15.67, 17.0, 20.0, 21.0, 21.5, 22.0, 22.5, 23.0, 23.25, 23.5, 23.75, 24.0, 24.25, 24.5, 24.75, 25.0, 25.25, 25.5, 26.0, 27.0
The range from 20 days to 26 days showed predictable variability for both noise values. (Noise (ppm) = 100 was also tested for the three duration values for $\alpha$=.3, $\alpha$=.5, and $\alpha$=.9 models) as shown in Figure 13 and 14 (figure 13 scaled up to 20 days to 27 days).
Figure 13: Recall vs Anomaly Rate across Duration values for Noise = 50, 225. Yellow line is for n_curves=300, blue for n_curves=100
Figure 14: Recall vs Anomaly Rate across Duration values for Noise = 50, 225. Yellow line is for n_curves=300, blue for n_curves=100
Final alpha variant model selections were done using this revised combined scoring metric:
$S_{\text{combined}}(\alpha)$ $=w_{\text{recall}}\cdot R_{\text{norm}}+w_{\text{anomaly}}^{-1}\cdot A_{\text{inv}}+w_{\delta}\cdot\Delta_{\text{norm}}$
$\quad+w_{\text{importance}}\cdot I_{\text{norm}}+w_{\text{range}}\cdot P_{\text{inv}}$ (7)
where the weight assignments follow an $\alpha$-controlled group structure:
Detection Group (weight = $\alpha$):
$w_{\text{recall}}$ $=0.5\alpha$ (8)
$w_{\text{anomaly}}^{-1}$ $=0.2\alpha$ (9)
$w_{\delta}$ $=0.3\alpha$ (10)
Interpretability Group (weight = $1-\alpha$):
$w_{\text{importance}}$ $=0.5(1-\alpha)$ (11)
$w_{\text{range}}$ $=0.5(1-\alpha)$ (12)
and the normalized/inverted metrics are defined as:
$R_{\text{norm}}$ $=\text{MinMaxScaler}(\text{TOI Recall})$ (13)
$A_{\text{inv}}$ $=1-\text{MinMaxScaler}(\text{Anomaly Rate})$ (14)
$\Delta_{\text{norm}}$ $=\text{MinMaxScaler}(\text{Recall-Anomaly Delta})$ (15)
$I_{\text{norm}}$ $=\text{MinMaxScaler}(\text{Mean Importance Top 20\%})$ (16)
$P_{\text{inv}}$ $=1-\text{MinMaxScaler}(\text{Top 10\% Range Upper})$ (17)
This revised scoring function retains convexity: all five component metrics are normalized to $[0,1]$ and weighted by non-negative coefficients summing to 1, yielding a bounded, interpretable score. The weights are structured such that the total contribution from the Detection group is controlled by $\alpha$, and the Interpretability group by $(1-\alpha)$, preserving the convex combination property and ensuring no extrapolation beyond the empirical performance ranges.
The revised scoring metrics is justified in terms of maximizing the weighted sum of five components as follows:
Detection Group ($\alpha$ weight):
  1. 1.
[…]
  2. 2.
[…]
  3. 3.
[…]
Interpretability Group ($1-\alpha$ weight):
  1. 1.
    Feature Importance (0.5($1-\alpha$)) - maximize clarity of top 20% features
  2. 2.
    Inverted Range Upper (0.5($1-\alpha$)) - minimize prediction instability in top 10%
A formal derivation of the combined scoring metric and its $\alpha$-controlled weighting scheme remains an open direction for future work as discussed under the banner of Theoretical Framework Development section of future works discussion (6.5).
Following are the urf-4 alpha variant models that were picked for the three alpha values:
  • $\alpha_{0.3}$ - urf4_model_variant_n300_d23.75_c2_n50ppm
    • Predicted: TOI Recall = 0.64; Anomaly Rate = 0.58; Mean Importance Top 20P = 0.42, Top 10Pimportance Upper Limit = 2.5
  • $\alpha_{0.5}$ - urf4_model_variant_n300_d25.25_c2_n225ppm
    • Predicted: TOI Recall = 0.73; Anomaly Rate = 0.62; Mean Importance Top 20P = 0.45, Top 10Pimportance Upper Limit = 19.7
  • $\alpha_{0.9}$ - urf4_model_variant_n300_d24.5_c2_n225ppm
    • Predicted: TOI Recall = 0.85; Anomaly Rate = 0.68; Mean Importance Top 20P = 0.37, Top 10Pimportance Upper Limit = 81.9
[…]
4.2 Weighted Root Sum Score: A revision of URF anomaly scores
As observed previously, the low $\alpha$ model is conservative over flagging TOIs and does better at scoring important curves higher. Meanwhile, the high $\alpha$ model is significantly more aggressive in flagging curves and casts a wider net with higher TOI recall. To allow ourselves hybrid control over aggressiveness and conservativeness, we needed a combined scoring function for the intersection set of the anomalies (non zero anomaly score curves) flagged by these two models. Thus, we pick the weighted root sum score (WRSS):
$s_{w}=w_{3}\,\sqrt{s_{3}^{\prime}}+w_{9}\,\sqrt{s_{9}^{\prime}}$ (18)
where $s_{3}^{\prime}$ and $s_{9}^{\prime}$ denote the raw anomaly scores from models $\alpha=0.3$ and $\alpha=0.9$ respectively, and $s_{w}$ is the weighted score computed from their square roots. The weights $w_{3}$ and $w_{9}$ satisfy:
$w_{3}+w_{9}=1$ (19)
To give higher preference to better scoring, we set the weights as $w_{3}=0.8$ and $w_{9}=0.2$. For a standardized comparison across urf-4 model runs on different test sets, we normalize $s_{w}$ (N-WRSS) using:
$\hat{s}_{w}=\frac{s_{w}-\min(s_{w})}{\max(s_{w})-\min(s_{w})}$ (20)
Figure 15: Distribution of N-WRSS g.t.e 0.4 on the intersection set of $\alpha=0.3$ and $\alpha=0.9$ models run on TESS SPOC Sector 1 Curves
Figure 16: Distribution of N-WRSS g.t.e 0.6 on the intersection set of $\alpha=0.3$ and $\alpha=0.9$ models run on TESS SPOC Sector 1-5 Curves
4.2.1 Why compute N-WRSS for $\alpha_{0.3}$ - $\alpha_{0.9}$ intersection sets?
The normalized weighted root sum of squares (N-WRSS $\hat{s}_{w}$) provides a joint metric that captures the agreement or divergence between different URF models ($\alpha_{0.3}$ and $\alpha_{0.9}$ in this case). By computing $\hat{s}_{w}$ over the intersection set - i.e., light curves flagged by both models - we:
  • Quantify inter-model consistency: A low $\hat{s}_{w}$ implies the two models assign similar anomaly scores (after normalization), reinforcing confidence in the flagged anomalies.
  • Suppress model-specific bias: Since $\alpha_{0.3}$ and $\alpha_{0.9}$ may differ in sensitivity due to their hyperparameter tuning or training noise characteristics, $\hat{s}_{w}$ acts as a joint view, minimizing the influence of any one model’s noise.
  • Support comparative benchmarking: It gives a unified score to evaluate intersection anomalies, useful when testing the consistency or precision of later similarity-based matching (e.g., cosine similarity labeling).
Thus, $\hat{s}_{w}$ on the $\alpha_{0.3}$ - $\alpha_{0.9}$ intersection forms a critical bridge for robust scoring before downstream similarity analysis or labeling.
In its current implementation, the N-WRSS reduces score variance and compresses outlier influence in a computationally tractable way. For further mathematical and theoretical justification of this, refer to the Theoretical Framework Development section of Section 6.5.
4.3 Astrometric and Stellar Parameter Augmentation
To enable physical interpretation of anomalies detected by URF-4, we enrich high-scoring candidates with astrophysical and astrometric metadata. Specifically, we extract relevant features from the TESS Input Catalog (TIC) and Gaia DR3 for all candidates exceeding a chosen anomaly threshold (in our case N-WRSS >=>= 0.4).
We begin by parsing TIC IDs from each candidate’s filename. These IDs are used to query the MAST TIC catalog to retrieve:
  • Astrometry: parallax (plx), proper motion in right ascension and declination (pmra, pmdec)
  • Stellar parameters: effective temperature (Teff), radius (rad), mass (mass), luminosity (lum), and surface gravity (logg)
To further characterize stellar kinematics, we compute the tangential velocity $v_{\text{tan}}$, defined as:
$v_{\text{tan}}=4.74\times\frac{\sqrt{\mu_{\text{RA}}^{2}+\mu_{\text{DEC}}^{2}}}{\pi}$
where $\mu_{\text{RA}},\mu_{\text{DEC}}$ are the proper motions in milliarcseconds per year, and $\pi$ is the parallax in milliarcseconds.
To assess the reliability of Gaia astrometric solutions, we query the Gaia DR3 catalog for the Renormalized Unit Weight Error (RUWE) of each source, a proxy for fit quality and potential binarity or extended structure.
The resulting dataset comprises both anomaly metadata and astrophysical properties, enabling cluster-level interpretation. The full set of appended features includes:
plx, pmra, pmdec, Teff, rad, mass, lum, logg, $v_{\text{tan}}$, ruwe
Refer to Appendix B for a descriptive list of these features.
4.4 Evaluating Astrometric Feature Discriminability
To assess which combinations of astrometric and stellar parameters most effectively distinguish high-anomaly candidates, we perform a t-SNE embedding followed by silhouette score ranking of feature sets on Sector 1 URF-4 $\alpha_{0.3}\cap\alpha_{0.9}$ set.
We generate all combinations of two or more features from the enriched feature set (Section 4.3), and for each subset:
  1. 1.
    Apply standard scaling to the selected features after removing NaN entries.
  2. 2.
    Use t-SNE to project the high-dimensional feature space into two dimensions.
  3. 3.
    Label each point as anomalous or not, based on a binary threshold (0.6 (arbitrary) for our case) on the URF-derived N-WRSS.
  4. 4.
    Compute the silhouette score for the binary labels in the embedded space.
The silhouette score provides a measure of cluster separation, with higher values indicating that the selected feature combination better separates anomalous from typical light curves in the reduced space.
This process identifies the most discriminative physical features for downstream interpretation, linking anomaly detection outcomes to astrophysical structure.
These findings provide an initial quantitative basis for linking astrometric and stellar parameters to anomaly clustering behavior. Future work (6.5) will expand this into a formal statistical validation framework, including comparison with established methods (e.g. Crake & Martínez-Galarza (2023); Melton et al. (2024)) and deeper analysis of why certain features like Teff, RUWE, and $v_{\mathrm{tan}}$ exhibit strong correlations with high N-WRSS anomalies, thereby enabling more grounded astrophysical interpretation of detected outliers.
4.5 Results
  • The top-performing set was: [’pmdec’, ’Teff’, ’vtan’, ’ruwe’] with a silhouette score of  0.069. Top 10 feature sets are provided in Table 6
  • The most visually separable and physically interpretable set was: [’Teff’, ’vtan’, ’ruwe’], showing excellent concentrated red region clusters.
  • Simpler sets such as [’Teff’, ’ruwe’] also performed well, confirming the significance of these features.
Table 6: Top 10 astrometric feature sets ranked by silhouette score based on their separation of high- and low-anomaly candidates in t-SNE space.
Rank Feature Set Silhouette Score
1 [pmdec, Teff, vtan, ruwe] 0.069
2 [Teff, ruwe] 0.068
3 [pmra, pmdec, Teff] 0.064
4 [pmra, rad] 0.062
5 [rad, vtan] 0.056
6 [Teff, rad] 0.055
7 [plx, pmdec, rad, logg] 0.053
8 [pmra, pmdec, rad, logg] 0.052
9 [plx, pmra, rad, lum] 0.051
10 [pmra, pmdec, Teff, vtan, ruwe] 0.051
Figure 17: t-SNE clustering of Sector 1 N-WRSS >= 0.6 curves based on [pmdec, Teff, vtan, ruwe]. N-WRSS goes from blue to red increasing from 0.0 to 0.9 (non-linearly to mark N-WRSS >= 0.6 better visually)
Figure 18: t-SNE clustering of Sector 1 N-WRSS >= 0.6 curves based on [Teff, vtan, ruwe]. N-WRSS goes from blue to red increasing from 0.0 to 0.9 (non-linearly to mark N-WRSS >= 0.6 better visually)
4.6 Next Steps: Cross-Sector Consistency and Physical Interpretation
To build on the silhouette-based feature evaluation, we propose a focused cross-sector analysis using two key feature sets: the top-scoring set [pmdec, Teff, vtan, ruwe] and the most visually separable set [Teff, vtan, ruwe]. We will apply these feature combinations to high-N-WRSS URF-4 $\alpha_{0.3}\cap\alpha_{0.9}$ candidates across four TESS sectors to evaluate the consistency of anomaly clustering in astrometric and stellar parameter space.
The objective is to determine whether URF-identified high scoring anomalies consistently occupy specific, distinguishable regions in this astrophysical feature space.
This multi-sector analysis will serve as an important validation step, linking unsupervised anomaly detection results to underlying physical features and assessing the reproducibility of clustering structures in high-dimensional parameter space. Results of this cross sector run are presented in figures 19 and 20.
4.7 Interpretation of Cross-Sector Embeddings
The t-SNE embeddings in Figures 19 and 20 reveal consistent structural patterns across TESS Sectors 1–4. In particular, the [Teff, vtan, ruwe] projection shows a recurring concentration of high N-WRSS scores (red regions) along narrow manifolds in the reduced space, suggesting the existence of a physically distinct anomaly subpopulation. The inclusion of proper motion in the extended feature set [pmdec, Teff, vtan, ruwe] enhances this separability, with clearer filamentary structures observed. This reinforces the hypothesis that high-scoring anomalies are not randomly distributed but instead correlate with specific stellar or kinematic regimes, validating the physical relevance of the URF-derived anomaly scores across sectors.
Figure 19: SECTOR 1 - 4 (CLOCKWISE FROM TOP LEFT) on feature set TEFF VTAN RUWE with N-WRSS on same color scale as Figure 19
Figure 20: SECTOR 1 - 4 (CLOCKWISE FROM TOP LEFT) on feature set TEFF VTAN RUWE with N-WRSS on same color scale as Figure 19
4.8 SIMBAD Weak Supervision and Morphological Similarity Matching
To bridge anomaly detection with astrophysical classification, we introduce a weakly supervised labeling approach using the SIMBAD object taxonomy. A subset of high-anomaly light curves (TOIs from Sector 2 - specific choice of sector in this case is arbitrary - for which Simbad Object Labels could be found) is cross-matched with MAST astronomical coordinates to retrieve SIMBAD labels such as Pl, EB*, V*, and others. This offers valuable physical priors without introducing strong biases. These labeled instances are used to construct a Sector 2 reference set of TOIs with known SIMBAD classifications, enabling downstream morphological matching and interpretation.
We extract feature vectors from these labeled TOIs using Lomb-Scargle periodograms and flux values of the curves as described in 3.1, apply PCA for dimensionality reduction, and save the compressed representations. These features form a labeled reference basis for similarity-based classification.
Subsequently, we compute pairwise cosine similarity between urf-4 $\alpha_{0.3}\cap\alpha_{0.9}$ set candidates and this labeled reference set. Each anomaly is assigned the closest-matching SIMBAD label group based on morphological similarity. Following label groups have been defined for this purpose:
    label_groups = {
    "planet_like": ["Pl", "Pl?", "BD*", "s*b"],
    "binary_star": ["SB*", "**", "EB*", "LM*"],
    "stellar": ["*", "PM*", "Er*"],
    "outlier": ["outlier", "err"]
The planet_like label group comprises synthetic light curves designed to exhibit transit-like morphology i.e., localized, symmetric flux dips with flat out-of-transit baselines. These shapes mimic the expected photometric signature of transits and are intended to guide anomaly detection models toward identifying signals with similar temporal structures. By concentrating on this subset, we enable both training and evaluation phases to focus on transit-relevant anomalies, filtering out unrelated variability patterns such as flares, rotational modulation, or systematics.
This step enables astrophysically interpretable annotation of previously unlabeled anomalies, allowing further insight into their likely nature and group clustering behavior. The methodology supports scalable, domain-aware classification while maintaining the unsupervised integrity of the original detection pipeline.
Figure 21: Distribution of N-WRSS from Sectors 1 - 5 by Matched Label from Cosine Similarity with a threshold of 0.6 for similarity score
4.8.1 Weak Label Acquisition via SIMBAD Cross-Matching
Since SIMBAD has scarce labeling when compared to the sheer number of tess curves, for a targeted and valuable labeled feature set, we chose to query the catalog for Object Type of TOIs from a specific sector. For the TIC identifier of each TOI, we queried a radius of 5 arcseconds around the right-ascension and declination for the TIC obtained from the MAST catalog. Table 10 lists the frequencies for each label for our reference feature set.
4.8.2 Morphological Classification via Cosine Similarity
With the Sector 2 TOIs now labeled and represented in a reduced-dimensional feature space (refer to Table 10 for similarity reference dataset label occurence frequencies), we proceed to classify unlabeled anomalies from Sectors 1 to 5 using morphological similarity. For each anomaly in the urf-4 $\alpha_{0.3}\cap\alpha_{0.9}$ set, we extract Lomb-Scargle and flux-based features, project them using the PCA model trained on the labeled reference set, and compute cosine similarity with all reference vectors.
A helper function match_fits_by_cosine_similarity performs the core matching step by:
  1. 1.
    Extracting feature vectors from the input light curve using the same pipeline used for the labeled TOIs as discussed in Section 3.1.
  2. 2.
    Projecting the feature vector into the same PCA space as the reference set.
  3. 3.
    Computing cosine similarity between the projected anomaly and all labeled reference TOIs.
  4. 4.
    Returning the highest similarity match above a threshold, or marking the sample as an “outlier” if no match exceeds the threshold.
To scale this process, we implement a batch version run_cosine_matching_batch that operates over the full anomaly candidate set using parallel workers. We apply a minimum similarity threshold (min_sim_threshold=0.6) to filter weak or ambiguous matches. The resulting matches are enriched with SIMBAD labels and their corresponding high-level label groups (as defined in label_groups) and merged back into the anomaly dataframe.
This classification approach allows for the astrophysically guided annotation of anomaly candidates, making it possible to investigate the clustering of morphologically similar signals and to validate the physical realism of urf-identified anomalies.
4.8.3 Results of Morphological Classification via Cosine Similarity
To assess the astrophysical plausibility of morphologically matched labels, we analyze the distribution of URF anomaly scores across the assigned SIMBAD label categories.
We visualize the results via strip plots of the N-WRSS against matched SIMBAD labels. To interpret these distributions, we define three separate diagnostic views:
  1. 1.
    Full distribution (Figure 21): All matched candidates, regardless of anomaly score, are shown.
  2. 2.
    High similarity & low anomaly score (Figure 22): We restrict to candidates with N-WRSS $\leq 0.3$ and cosine similarity $\geq 0.9$.
  3. 3.
    High anomaly score & low similarity (Figure 23): This diagnostic isolates candidates with anomaly scores $\geq 0.85$ and cosine similarity $\leq 0.3$.
Figure 22: Distribution of Low N-WRSS and High Similarity Score from Sectors 1 - 5 by Matched Label from Cosine Similarity with a threshold of 0.6 for similarity score.
Figure 23: Distribution of High N-WRSS and Low Similarity Score from Sectors 1 - 5 by Matched Label from Cosine Similarity with a threshold of 0.6 for similarity score (shows no curves with high anomaly score that could not be successfully classified into one of the SIMBAD Object type based morphology labels for the light curves)
4.8.4 Interpretation of Cosine Similarity Results
The three diagnostic views (figures 21, 22 and 23) enable a structured assessment of how anomaly scores and morphological similarity align with known astrophysical classes.
First, the full distribution (Figure 21) confirms that the majority of anomalies are morphologically consistent with stellar variability or planetary candidates - with concentrated clusters in PM*, Pl, and * label groups and a significant population, with almost constant density across N-WRSS scores, of Planet-like label group above a threshold of N-WRSS=0.6 where other label groups have a much lesser and more scarce population.
Second, the high-similarity/low-anomaly set (Figure 22) demonstrates that URF correctly assigns low anomaly scores to morphologically standard light curves which can be classified into a labeled category. This supports the argument that low URF scores represent "normal" behavior in the training distribution.
Finally, and most critically, the high-anomaly/low-similarity view (Figure 23) yields zero candidates. This strongly indicates that nearly all high-score candidates still retain morphological similarity to known astrophysical classes - and that the pipeline’s anomaly scores are physically meaningful rather than driven by noise or artifacts.
Taken together, these plots show that cosine similarity provides astrophysically grounded, interpretable labels to anomaly candidates - and that URF scores correlate well with label group structure in a physically plausible way.
4.8.5 Cosine Similarity Label Group Statistics
In total, we processed 87,858 light curves spanning TESS SPOC Sectors 1 through 5. This forms the complete set of candidates from which anomaly scores were computed, morphological features extracted, and cosine similarity-based label assignments performed.
To evaluate the distribution of anomaly strength across label groups, we stratify all matched candidates by their assigned label group (planet_like, stellar, binary_star) and analyze how the number of candidates varies with increasing N-WRSS (normalized weighted root sum score) threshold. The plots in figures 24, 25 and 26, for each label group, show the number of candidates with anomaly score above a given threshold.
Figure 24: Planet-like candidates above n-wrss threshold for sectors 1 to 5
Figure 25: Stellar candidates above n-wrss threshold for sectors 1 to 5
Figure 26: Binary-star candidates above n-wrss threshold for sectors 1 to 5
Figure 27: First curve in cluster 1 of dbscan clustering on dtw matrix with eps 4.5 and min_samples 2
Figure 28: Second curve in cluster 1 of dbscan clustering on dtw matrix with eps 4.5 and min_samples 2
Figure 29: Silhouette score heatmap for grid search over eps and min_sample values. Two optimal values are eps = 2.0, min_samples = 10 and eps = 4.5, min_samples = 2
4.9 Clustering and Morphological Filtering Methods
To identify transit-like morphologies, we selected the planet_like classified group of curves from TESS Sectors 1 through 5. We applied a threshold on the normalized weighted root sum square (N-WRSS) anomaly score, retaining only candidates with N-WRSS $\geq 0.35$ (arbitrary threshold based on figure 21), resulting in a filtered set of 620 planet-like light curves for morphological clustering.
As an initial clustering approach, we employed DBSCAN using a distance matrix derived from Dynamic Time Warping (DTW) Salvador & Chan (2007), a technique well-suited for comparing time-series data with temporal misalignments or non-uniform sampling - conditions typical of TESS light curves.
To visualize the results, all curves were phase-folded using the most significant period from their Lomb-Scargle periodograms. A minimum power threshold was enforced to filter out curves with weak or unreliable periodicity.
We then performed enhanced morphological feature extraction on the remaining subset, replacing the feature set described in Section 3.1 to a refined 10-dimensional representation designed to capture morphological characteristics more effectively.
Multiple clustering algorithms were evaluated on this refined feature space, and their performance was assessed both quantitatively (via TOI membership tracking) and qualitatively (via visual inspection of folded curves within each cluster). The most effective clustering approach was selected based on its ability to group plausible transit-like signals and isolate visually coherent morphology types.
The results of these clustering experiments are presented thereafter.
Figure 30: 3 distinct morphological classes (each row represents a class) that populate the entirety of 0th cluster from DBSCAN clustering of DTW matrix
4.9.1 Dynamic Time Warping (DTW) and DBSCAN
For the 620 planet-like candidates, we first applied phase folding using the peak period from each curve’s Lomb-Scargle periodogram. The minimum power threshold filtered out the curves that did not have a Lomb Scargle period frequency with power greater than the threshold. The folding was performed using the following function from our pipeline:
clara2.phase_fold(fits_fpath,
                  min_period=0.2,
                  max_period=20,
                  ls_power_thresh=0.05)
This function returns the following key outputs:
  1. 1.
    Phase: Folded phase values of the light curve, sorted, ranging from $[0,1]$.
  2. 2.
    Folded flux: Flux values corresponding to each phase value.
  3. 3.
    Bin centers: Centers of the uniform phase bins used for averaging (default: 100 bins).
  4. 4.
    Bin means: Mean flux values per bin, representing the phase-folded light curve morphology.
To represent each curve’s morphology independently of sampling density, we used the bin means as fixed-length feature vectors. These vectors were padded or truncated to a standard length, normalized, and stored in a list called features. We then computed the Dynamic Time Warping (DTW) pairwise distance matrix using the dtaidistance library:
n = len(features)
D = np.zeros((n, n))
# Compute pairwise DTW distances
for i in range(n):
    for j in range(i+1, n):
        d = dtw.distance_fast(features[i], features[j])
        D[i, j] = D[j, i] = d
With the symmetric distance matrix $D$ computed, we applied DBSCAN for clustering. DBSCAN has two primary hyperparameters:
  1. 1.
    eps: Maximum distance between two points to be considered neighbors.
  2. 2.
    min_samples: Minimum number of neighbors required to form a core point.
To identify optimal clustering parameters, we performed a grid search over combinations of eps and min_samples, evaluating each using the silhouette score. The results are shown in Figure 29.
The two best configurations from the grid search were:
  1. 1.
    $\epsilon=2.0$, min_samples = 10 → 2 clusters, 203 noise points
  2. 2.
    $\epsilon=4.5$, min_samples = 2 → 2 clusters, 29 noise points
Since the DTW-based clustering step is used as a pre-filtering mechanism for removing outliers prior to more refined morphological clustering, we selected the second configuration ($\epsilon=4.5$, min_samples = 2), which yielded a cleaner separation with fewer noise points.
This configuration produced:
  • 29 curves in the DBSCAN noise cluster (1-1),
  • 453 curves in cluster 0,
  • 2 curves in cluster 1.
The two curves in cluster 1 are presented in Figures 27 and 28. Cluster 0, which contained the 453 curves (filtered down from the initial 620 curves), exhibited three distinct morphological types, shown in Figure 30. This cluster also contained 19 TOIs. Neither cluster 1-1 nor cluster $1$ contained any TOIs.
4.9.2 Feature of a Curve - II
For improved clustering and morphological separation of the visually identified classes in cluster 0 (from DBSCAN applied on DTW distances), we derived a 10-dimensional feature vector for each candidate curve. This initial 10-dimensional representation serves as a baseline for morphological analysis and transit-like anomaly clustering. Future work (see Section 6.5 under Advanced Morphological Analysis) will extend this by exploring more expressive feature combinations based on these and more. These 10 features are as follows:
  1. 1.
    Transit Depth is defined as the difference between the minimum and median of the binned folded flux:
    $d=F_{\text{min}}-F_{\text{median}}$ (21)
    where:
    $F_{\text{min}}$ $=\min_{i}F_{i}$
    $F_{\text{median}}$ $=\text{median}_{i}(F_{i})$
    and $F_{i}$ denotes the binned phase-folded flux.
    Simpler explanation: How deep the transit dip is compared to the average light level.
  2. 2.
    Transit Width: Given the folded light curve phase values $\{\phi_{i}\}$ and corresponding binned flux $\{F_{i}\}$:
    $F_{\text{min}}$ $=\min_{i}F_{i}$ (22)
    $F_{\text{max}}$ $=\text{median}_{i}(F_{i})$ (23)
    $F_{\text{half}}$ $=\frac{F_{\text{min}}+F_{\text{max}}}{2}$ (24)
    Let $\mathcal{I}=\{i\mid F_{i}<F_{\text{half}}\}$ be the set of indices below half-depth. Then:
    $w=\begin{cases}\phi_{\max(\mathcal{I})}-\phi_{\min(\mathcal{I})},&\text{if }|\mathcal{I}|\geq 2\\ 0,&\text{otherwise}\end{cases}$ (25)
    Simpler explanation: The width of the transit in phase space, indicating how long the dip lasts.
  3. 3.
    Baseline Standard Deviation is computed from the top 20% of folded flux values (assumed out-of-transit):
    $\mathcal{B}$ $=\left\{F_{i}\mid F_{i}>P_{80}(F)\right\}$ (26)
    $\sigma_{\text{baseline}}$ $=\text{std}(\mathcal{B})$ (27)
    Simpler explanation: Measures the flux variation outside the transit, acting as a noise proxy.
  4. 4.
    Asymmetry Index quantifies flux asymmetry around the center of the phase curve:
    $N$ $=\text{length of }F$ (28)
    $M$ $=\left\lfloor\frac{N}{2}\right\rfloor$ (29)
    $a$ $=\frac{1}{M}\sum_{i=1}^{M}\left|F_{i}-F_{N-i+1}\right|$ (30)
    Simpler explanation: Indicates whether the transit shape is symmetric or skewed.
  5. 5.
    Sharpness is the discrete curvature at the flux minimum after smoothing:
    $\tilde{F}$ $=\text{GaussianSmooth}(F,\sigma=3)$ (31)
    $i_{\min}$ $=\arg\min_{i}\tilde{F}_{i}$ (32)
    $s$ $=\begin{cases}\tilde{F}_{i_{\min}-1}-2\tilde{F}_{i_{\min}}+\tilde{F}_{i_{\min}+1},&\text{if }2<i_{\min}<N-3\\ 0,&\text{otherwise}\end{cases}$ (33)
    Simpler explanation: Measures how "pointed" or sharp the dip is at its lowest point.
  6. 6.
    Autocorrelation Strength is the highest non-zero autocorrelation peak:
    $\bar{F}$ $=\frac{F-\mu}{\sigma}$ (34)
    $\text{ACF}_{k}$ $=\sum_{i=1}^{N-k}\bar{F}_{i}\cdot\bar{F}_{i+k}$ (35)
    $\text{ACF}_{0}$ $=0$ (36)
    $a$ $=\max\left\{\text{ACF}_{k}\mid k>0\text{ and }\text{ACF}_{k}\text{ is a local maximum}\right\}$ (37)
    Simpler explanation: Quantifies repeating structure or periodicity in the shape of the curve.
  7. 7.
    Transit Count is the number of strong peaks in the Lomb-Scargle periodogram:
    $F_{\text{norm}}$ $=\frac{F-\text{median}(F)}{\text{std}(F)}$ (38)
    $\text{LS}(\omega)$ $=\text{LombScargle}(t,F_{\text{norm}})$ (39)
    $P(\omega)$ $=\text{LS.power}(\omega)$ (40)
    $\mathcal{T}$ $=\left\{\omega_{i}\mid P(\omega_{i})\text{ is a peak with prominence }>\delta\right\}$ (41)
    Transit Count $=|\mathcal{T}|$ (42)
    Simpler explanation: Counts how many strong periodic signals are present.
    We use the Box Least Squares (BLS) algorithm to identify periodic, box-shaped dips in normalized light curves—an effective method for detecting transiting exoplanets. BLS iteratively searches over trial periods to find the one that maximizes a transit signal detection statistic, fitting a simple box model to approximate the flux drop caused by a planet transit. From the best-fit period $P^{\ast}$, we derive the last three of the ten interpretable metrics that characterize the transit-like signal:
  8. 8.
    BLS Transit Depth is the best-fit transit depth from a box-shaped model:
    $d_{\text{BLS}}=\text{depth}[P^{\ast}]$ (43)
    where $P^{\ast}=\arg\max_{P}\text{BLS}_{\text{power}}(P)$.
    Simpler explanation: Best-estimate of how deep the box-like transit is.
  9. 9.
    BLS Transit Duration is the duration of the best-fit box-shaped transit:
    $T_{\text{dur}}=\text{duration}[P^{\ast}]$ (44)
    Simpler explanation: How long the transit lasts in time, according to the box model.
  10. 10.
    BLS Transit Signal-to-Noise Ratio (SNR):
    $\text{SNR}_{\text{BLS}}=\frac{d_{\text{BLS}}}{\sigma_{F}}$ (45)
    where $\sigma_{F}=\text{std}(F_{\text{norm}})$.
    Simpler explanation: Strength of the detected box transit signal relative to background noise.
Table 7: Comparison of clustering methods applied to the 10-dimensional feature set extracted from URF-4 planet-like candidates. Each entry lists the number of TOIs, total curves, and the percentage of TOIs in that cluster.
Clustering Method Cluster Label TOIs Curves Percentage (%)
HDBSCAN Cluster -1 3 155 1.9
Cluster 0 16 263 6.1
Cluster 1 0 35 0.0
KMeans (k=3) Cluster 0 11 228 4.8
Cluster 1 1 144 0.7
Cluster 2 7 81 8.6
DBSCAN Cluster -1 2 68 2.9
Cluster 0 16 304 5.3
Cluster 1 1 57 1.8
Cluster 2 0 5 0.0
Cluster 3 0 7 0.0
Cluster 4 0 5 0.0
Cluster 5 0 4 0.0
Cluster 6 0 3 0.0
Bayesian GMM (n=2) Cluster 0 16 174 9.2
Cluster 1 3 279 1.1
GMM (n=2) Cluster 0 17 155 11.0
Cluster 1 2 298 0.7
DPMM (n=3) Cluster 0 2 175 1.1
Cluster 1 1 164 0.6
Cluster 2 16 114 14.0
4.9.3 Comparison and Interpretation of Clustering Outcomes
To explore the morphological diversity among the 620 planet-like candidates, we applied a range of clustering algorithms to a 10-dimensional feature set engineered from phase-folded light curves. This feature set is as described in section 4.9.2. Our objective was to determine whether these features could reveal structural subgroups and isolate known TOIs within well-defined clusters. Results of clustering methods are shown in Table 7
Among all tested methods, Gaussian Mixture Models (GMM) with two components produced the clearest separation: one cluster contained 17 TOIs among 155 curves, achieving the highest TOI density. Dirichlet Process GMM (DPMM) performed comparably well, grouping 16 TOIs within a smaller cluster of 114 curves, indicating a stronger enrichment but with more fragmentation overall.
While other methods such as DBSCAN, HDBSCAN, and KMeans provided partial clustering success, they either included fewer TOIs or distributed them across broader clusters. The probabilistic nature of GMM and DPMM allowed them to better model the subtle variations in transit morphology, making them more suitable for this task.
In conclusion, the combination of interpretable features and flexible model-based clustering yielded the best separation of high-quality transit-like signals. The GMM and DPMM results (sampled in figures 34, 31, 32, and 33 to visualize morphological similarity) represent the most promising configurations to guide further candidate filtering, validation, and prioritization.
Figure 31: Random sample of curves from dpmm cluster 2
Figure 32: Random sample of curves from gmm cluster 0
Figure 33: Random sample from gmm cluster 1
Figure 34: Random sample from dpmm cluster 0
5 Computational Resources and Runtime
All computations were performed on a personal desktop running Ubuntu 24.04 LTS (kernel 6.11), equipped with an Intel® Core™ i3-8100 CPU (4 physical cores, 3.60 GHz), 32 GB of DDR4 RAM, and no discrete GPU. The system used integrated Intel UHD Graphics 630 for all display tasks. CLARA was developed and executed entirely on this CPU-based machine without GPU acceleration.
All code was written and executed in Jupyter notebooks launched using python3 -m notebook from within a Python virtual environment. The environment was configured with Python 3.12 and included standard scientific and astronomical computing libraries such as numpy, pandas, astropy, scikit-learn, tqdm, and batman-package.
Parallel processing was enabled via Python’s multiprocessing module. All multi-threaded tasks - including light curve downloading, URF scoring, and cosine similarity classification - were executed using up to 4 parallel workers, matching the CPU’s physical core count.
Downloading $\sim$16,000 2-minute cadence SPOC light curves from a single TESS sector took approximately 18 minutes using 4 threads. Running the full anomaly scoring phase (URF-4) on the same sector took $\sim$1.5 hours per $\alpha$ model (e.g., $\alpha=0.3$ scored 15,878 light curves in 89 minutes; $\alpha=0.9$ in 86 minutes). Cosine similarity classification across $\sim$7,400 light curves took approximately 34 minutes using 4 workers.
Additional preprocessing steps such as Lomb-Scargle computation, phase folding, and SIMBAD-based label similarity required $\sim$6 minutes for every 1,100 folded light curves. All tasks were executed on local storage spanning 3.7 TiB of disk capacity (including two external USB drives), with approximately 1.8 TiB utilized over the course of experimentation.
6 Discussion
6.1 Synthesis of Key Findings
The CLARA framework demonstrates that synthetic set design can systematically control the behavior of Unsupervised Random Forests for astronomical anomaly detection, addressing two fundamental challenges in large-scale time-series surveys: the need for interpretable, steerable and computationally efficient discovery mechanisms and the validation of detected anomalies against genuine astrophysical phenomena.
Our investigation reveals several critical insights into the relationship between synthetic training parameters and model performance. The systematic parameter exploration across 36 URF-4 subvariants establishes that given our six performance metrics, transit duration and noise level emerge as the most influential factors for steering model behavior, with transit duration dominating both TOI Recall and Anomaly Rate metrics while noise level primarily affects importance-based prioritization. This finding has immediate practical implications: researchers can now predictably tune anomaly detection systems toward specific scientific objectives by adjusting these key parameters within viable operational ranges (50-300 ppm for noise, 13-27 days for duration). More finer mapping of feature to performance metric using the a suitable combined scoring metric using a single $\alpha$ coefficient or multiple coefficients can be done to predict urf behavior better.
The $\alpha$-parameterized scoring framework successfully enables targeted control over the scientific trade-off between detection completeness and astrophysical significance. Cross-sector validation across TESS SPOC Sectors 1-5 confirms that this behavioral steering mechanism generalizes robustly, with $\alpha$=0.9 models consistently maximizing binary TOI recall (64% approx) and $\alpha$=0.3 models concentrating high-value targets within narrow score ranges (1.6-9.7% percentile span (approx)). This consistency across diverse stellar populations validates the transferability of synthetic training approaches and enables confident deployment without sector-specific retraining, especially when a combined anomaly scoring metric like N-WRSS is used.
6.2 Astrophysical Validation and Physical Interpretability
The correlation analysis between anomaly scores and stellar parameters provides compelling evidence that URF-detected anomalies correspond to genuine astrophysical phenomena rather than instrumental artifacts. The t-SNE embedding analysis using astrophysical features [Teff, vtan, ruwe] reveals that high-scoring anomalies cluster along distinct manifolds in parameter space, suggesting they represent a physically coherent subpopulation rather than random noise.
Particularly significant is the morphological classification validation through SIMBAD cross-matching and cosine similarity analysis. The complete absence of high-anomaly candidates (N-WRSS $\geq$ 0.85) with low morphological similarity (cosine similarity $\leq$ 0.3) to known astrophysical classes strongly supports the hypothesis that URF anomaly scores reflect genuine astrophysical structure. This finding addresses a fundamental concern in unsupervised anomaly detection: whether statistical "weirdness" corresponds to scientifically meaningful phenomena. The consistent morphological similarity of high-scoring candidates to established planetary, stellar, and binary star classes validates the physical realism of the detection framework.
The Normalized Weighted Root Sum of Squares (N-WRSS) scoring scheme effectively combines the complementary strengths of conservative ($\alpha$=0.3) and aggressive ($\alpha$=0.9) models, providing a unified metric that captures both consensus-based confidence and comprehensive coverage. The distribution analysis across SIMBAD object classes reveals that planet-like candidates maintain higher representation at elevated N-WRSS thresholds compared to stellar or binary classifications, suggesting that the framework exhibits appropriate sensitivity to transit-like morphologies.
6.3 Computational Efficiency and Methodological Advances
The CLARA framework introduces several methodological innovations that advance the state of unsupervised astronomical anomaly detection. The systematic exploration of synthetic parameter space moves beyond the fixed hyperparameter approaches of previous work, enabling researchers to understand and control the mapping between input design choices and detection outcomes. This controllability represents a paradigm shift from black-box anomaly detection toward interpretable, goal-oriented discovery systems.
The computational efficiency achieved on modest hardware (4-core i3 CPU, 32GB RAM) demonstrates the practical scalability of the approach. Processing 1̃6,000 to 20,000 light curves per sector in approximately 80 minutes to 120 minutes per $\alpha$ model (it varies on number of curves and value of $\alpha$ - higher alpha models take lesser time than lower alpha models which behave more conservatively), with full anomaly scoring, feature extraction, and morphological classification completing within reasonable timeframes, validates the framework’s suitability for large-scale surveys.
The modular pipeline design facilitates extensibility to other survey architectures and scientific objectives. The clear separation between synthetic set generation, anomaly scoring, and astrophysical validation enables researchers to adapt individual components while preserving the overall framework integrity. This modularity is particularly valuable for extending CLARA to Full Frame Image (FFI) data, where the broader stellar population and different noise characteristics would require adjusted synthetic training strategies while maintaining the core methodological approach.
6.4 Limitations and Scope of Applicability
Several important limitations constrain the generalization of current results. The restriction to 2-minute cadence SPOC targets introduces selection bias toward brighter, pre-selected stars, potentially limiting applicability to the broader TESS population accessible through FFI data. Our analysis suggests that FFI-based light curves would establish different baselines for "normal" behavior, requiring recalibration of anomaly score thresholds and potentially altering the specific mappings between synthetic parameters and performance metrics. While the core hypothesis, that synthetic set design steers model behavior, should remain valid, the quantitative relationships established here may not transfer directly to FFI populations.
The evaluation framework relies heavily on TOIs as ground-truth proxies, which introduces potential circularity in performance assessment. TOIs themselves represent a biased sample, selected through existing detection algorithms and subject to observational and instrumental selection effects. This dependence limits our ability to assess the discovery of genuinely novel phenomena that fall outside current detection paradigms. The morphological classification via SIMBAD cross-matching partially addresses this limitation by providing independent astrophysical validation, but the sparse labeling in SIMBAD restricts the scope of this validation approach.
The synthetic light curve generation employs simplified transit models (box-shaped and Mandel-Agol) with uniform parameter distributions, which may not adequately represent the full diversity of real planetary systems. More sophisticated synthetic generation incorporating realistic parameter correlations, stellar activity, and instrumental systematics, if possible to be tested under controlled changes in the parameter values, could improve model performance and generalization. Similarly, the feature extraction approach, while computationally efficient, truncates light curves to 3000 flux points and may miss long-duration or late-occurring anomalies.
6.5 Future Directions and Extensions
The CLARA framework establishes a foundation for several promising research directions. Based on the comprehensive table of planned improvements (see author’s development notes), we outline nine key areas for future investigation:
Immediate Technical Extensions: The framework will be scaled to a proposed 500,000 light curve benchmark. This will test the scalability of the CLARA pipeline at clara-cli.
Enhanced Statistical Validation: Future work will implement rigorous statistical validation frameworks to quantify detection significance and establish confidence intervals for anomaly classifications. This includes comparison studies with established methods like Crake & Martínez-Galarza (2023), Melton et al. (2024) etc. incorporating their insights. Additionally, correlation analysis between stellar parameters (Teff, RUWE, vtan) and N-WRSS scores, more specifically, quantifying why the stellar parameters that correlate well with better high N-WRSS clustering tend to do so, will provide deeper astrophysical interpretation of detected anomalies.
Advanced Morphological Analysis: The morphological clustering analysis using engineered features represents an initial step toward automated anomaly taxonomy. Planned expansions include: (1) sophisticated feature combination studies beyond the current 10-dimensional representation as in section 4.9.2, (2) integration of deep learning-based representations for more nuanced morphological classification, and (3) development of improved metrics that surpass current TOI-based and other metrics as in Table 3. The combination of unsupervised detection with supervised morphological classification offers a promising hybrid approach for systematic discovery of new astrophysical phenomena.
Cross-Survey Applications: Extension to other time-series surveys (Kepler, LSST, K2) will test the broader applicability of controllable synthetic training approaches. Each survey presents unique challenges in sampling synthetic set features, and astronomical populations, requiring adapted synthetic generation strategies while preserving core framework principles. The modular design of CLARA facilitates such extensions by isolating survey-specific components from the general methodological approach.
Theoretical Framework Development: A key future priority is to formalize the theoretical foundations of the Normalized Weighted Root Sum Score (N-WRSS), which was introduced as a composite anomaly metric combining outputs from independently trained URF models. In its current formulation, weights are applied to the square roots of raw anomaly scores prior to normalization, primarily to reduce score variance and compress outlier influence in a computationally tractable way. However, this approach remains heuristic. A more rigorous theoretical analysis is needed to justify the use of square-root transformations, evaluate the statistical properties of such weighted combinations, and explore alternative aggregation strategies (e.g., geometric means, entropy-based weights, or learned fusion mechanisms).
In parallel, the revised performance scoring function used for alpha-controlled model selection, defined as a convex combination of normalized recall, anomaly suppression, and interpretability metrics, also warrants formal study. Although the current structure guarantees convexity and boundedness, its empirical weighting scheme (based on an $\alpha$-partitioned group structure) lacks statistical justification. Future work will explore principled approaches for assigning metric weights, sensitivity analysis across parameter regimes, and potential use of Pareto front optimization to balance competing objectives in a non-scalar fashion. Establishing rigorous foundations for both anomaly scoring and performance evaluation would enhance generalization beyond specific URF configurations and improve the interpretability and reliability of CLARA’s outputs across diverse stellar populations.
Automated Optimization: Development of meta-optimization deep learning based techniques for automatically learning optimal synthetic parameters represents a natural evolution beyond the current grid-search approach. Machine learning methods for optimizing synthetic set design based on target performance metrics could eliminate manual synthetic set generation parameter tuning and enable adaptive optimization for evolving scientific priorities. This is particularly valuable for multi-objective optimization scenarios where researchers must balance competing detection goals.
GAN-Driven Post-Filtering: A potential extension of CLARA involves training a post-clustering classifier to distinguish genuine transit-like curves from false positives. To avoid reliance on labeled TOIs, this model could be trained using GAN-generated light curves, where positive samples include synthetic Mandel-Agol transits injected into GAN-simulated stellar variability, and negatives consist of non-transit or eclipsing binary morphologies. The resulting classifier (e.g., CNN or contrastive encoder) would serve as a final, learnable filter after URF-based anomaly detection and morphological similarity based clustering (Cosine similarity, DTW, DBSCAN etc.), improving robustness against spurious signals.
6.6 Broader Context and Significance
The CLARA framework contributes to the broader evolution of astronomical methodology toward intelligent, unsupervised discovery systems capable of handling the data volumes generated by modern surveys. By demonstrating that unsupervised anomaly detection can be made both controllable and interpretable, this work addresses fundamental challenges in the transition from human-expert-driven discovery to machine-assisted exploration of astronomical parameter space.
The validation of detected anomalies through independent astrophysical catalogs establishes a methodology for bridging statistical pattern recognition with physical understanding. This integration is crucial for maintaining scientific rigor while leveraging the pattern detection capabilities of machine learning approaches. The framework provides a model for how automated discovery systems can incorporate domain knowledge and external validation to ensure astrophysical relevance of detected patterns.
The successful demonstration of behavioral steering through synthetic training represents a significant advance in making machine learning systems more transparent and controllable for scientific applications. Rather than treating unsupervised anomaly detection as a black-box process, CLARA enables researchers to understand and predict how methodological choices affect discovery outcomes. This transparency is essential for building confidence in automated unsupervised systems and ensuring that scientific priorities drive technological implementation rather than being constrained by algorithmic limitations.
In conclusion, the CLARA framework establishes both methodological foundations and practical tools for controllable, interpretable anomaly detection in large-scale astronomical surveys. The demonstrated ability to systematically steer detection behavior while maintaining astrophysical validity provides a pathway toward more intelligent, goal-oriented discovery systems capable of adapting to evolving scientific priorities and survey characteristics. As astronomical data volumes continue to grow exponentially, such frameworks will become increasingly essential for realizing the full scientific potential of modern survey missions.
7 Conclusions
We have presented CLARA, a modular and controllable framework for unsupervised transit detection using TESS light curves. Our two-part investigation demonstrates that synthetic training set design plays a critical role in shaping the behavior and effectiveness of Unsupervised Random Forest (URF) models for anomaly detection.
In Part I, we introduced URF-4, a synthetic-injection-based URF variant with tunable parameters, and performed a systematic parameter sweep across multiple synthetic configurations. By defining $\alpha$-dependent scoring metrics and balancing recall with astrophysical importance, we showed that synthetic training sets can be designed to prioritize either detection completeness ($\alpha=0.9$), astrophysical significance ($\alpha=0.3$), or a balanced trade-off ($\alpha=0.5$). We validated the generalization of these models across TESS Sectors 1–5 and demonstrated that synthetic design enables consistent and predictable performance aligned with distinct scientific goals. Our cross-sector generalization tests involved scoring 384,000 light curves (128,000 per $\alpha$-variant) sampled from five distinct TESS sectors, validating CLARA’s consistency across diverse stellar populations and observational conditions.
In Part II, we assessed the astrophysical relevance of URF-identified anomalies using stellar parameters from Gaia DR3, t-SNE cluster embeddings, and morphological similarity matching with SIMBAD-labeled TOIs. We introduced the Normalized Weighted Root Sum Score (N-WRSS) as a hybrid metric for reconciling high- and low-$\alpha$ models. This analysis revealed that high-scoring anomalies are concentrated in distinct astrophysical subspaces (e.g., in $\mathrm{T_{eff}}$$\mathrm{RUWE}$$\mathrm{v_{tan}}$ space), and exhibit morphological similarity to known exoplanetary and stellar variability classes.
Cosine similarity-based weak labeling showed that a majority of high-scoring anomalies align with planet-like or stellar variability categories, while no high-anomaly, low-similarity candidates were found - indicating the physical plausibility of detected anomalies and minimizing the risk of false positives due to noise.
Together, these results validate CLARA as a physically grounded, interpretable, and generalizable anomaly detection pipeline. By combining synthetic data-driven tuning, astrophysical correlation analysis, and unsupervised clustering techniques, CLARA offers a powerful toolkit for large-scale, label-free exploration of transit-like signals in time-domain surveys. Future work will extend this framework to include radial velocity cross-matching and application to other survey datasets beyond TESS.
Acknowledgements
This research was conducted entirely independently and without institutional support or funding. All computations were performed on a single personal desktop machine (Intel i3 processor, 32 GB RAM), which, despite its limitations, proved remarkably persistent. The author is deeply grateful for its service.
The author thanks the developers and maintainers of the open-source software and data archives acknowledged in the Data Availability section, which made this work possible.
Data Availability
The full CLARA pipeline, including Jupyter notebooks, source code, pre-trained URF models, and helper scripts for light curve preprocessing, anomaly scoring, and similarity matching, and the CLARA cli tool to execute the pipeline with the computational resources and efficiency as stated in section 5, is publicly available at: https://github.com/googleboy-byte/CLARA and https://github.com/googleboy-byte/clara-cli.
Intermediate results, including anomaly scores, N-WRSS scores, cosine similarity matches, and t-SNE cluster outputs, are provided as supplementary CSVs and figure files in the repository’s data/ and results/ directories.
All TESS light curve data used in this study were obtained from the Mikulski Archive for Space Telescopes (MAST) at https://archive.stsci.edu/tess/. Specific sectors analyzed include SPOC 2-minute cadence data from TESS Sectors 1–5.
TOI (TESS Object of Interest) catalogs were retrieved from the Exoplanet Follow-up Observing Program for TESS (ExoFOP-TESS): https://exofop.ipac.caltech.edu/tess/download_toi.php?sort=toi&output=csv
Stellar and astrometric parameters were retrieved from the TESS Input Catalog (TIC v8.2) and Gaia Data Release 3 (Gaia DR3), both accessible via the MAST and ESA Gaia Archive respectively.
Labeled reference sets for cosine similarity matching were constructed from TOIs cross-matched with SIMBAD object classifications. SIMBAD data were obtained from the Centre de Données astronomiques de Strasbourg (CDS): https://simbad.u-strasbg.fr/simbad/.
Synthetic light curves were generated using the batman transit model package (https://github.com/lkreidberg/batman). All synthetic parameter configurations, injection scripts, and random seeds used in this work are included in the GitHub repository to support full reproducibility.
Questions regarding the dataset or reproducibility materials should be directed to the corresponding author via email or github issues.
Appendix A Notes
A.1 URF-3 Cross-Sector Validation
URF-3’s performance exhibits significant sector-dependent variation due to its training methodology. Since URF-3 uses feature vectors extracted from confirmed TOIs in Sector 1 as its synthetic anomaly class, it demonstrates clear overfitting when evaluated on the same sector. This is evident in the score distribution plots (Figures 2 and 3), where URF-3’s anomaly scores for TOIs in Sector 1 are systematically higher (peaking around 0.45-0.47) compared to Sector 2 (peaking around 0.10-0.12). The reduced score separation in Sector 2 reflects URF-3’s diminished ability to generalize beyond the specific feature characteristics of its training TOIs. This sector dependency underscores the importance of independent validation sets and highlights a fundamental limitation of using confirmed target features directly as synthetic training data, reinforcing URF-4’s advantage in using parametric synthetic generation for improved generalization.
A.2 Hyper-parameter Search Space
n_estimators: 10 evenly spaced values between 50 and 200
max_features: [sqrt, log2]
max_depth: [100, 300, 500, 700, 900, 1000, None]
min_samples_split: [2, 4, 7, 10]
min_samples_leaf: [1, 2]
bootstrap: [True, False]
warm_start: [True, False]
A.3 Note on Duration Units
During the course of this research, the duration of transit-like events is consistently measured in days across all computations, analyses, and visualizations. However, in a few instances, particularly in intermediate plots or variable names, the duration may be mistakenly labeled as duration_hr. This is a human oversight in labeling only. All such instances refer to durations that are in fact expressed in days, not hours.
This clarification ensures consistency in interpretation and should be kept in mind when reviewing figures, tables, or variable definitions that include this mislabeled term.
A.4 Note on TOI Importance AUC
While TOI Importance AUC varied enough in value to give us an opportunity to analyze trends of synthetic parameters v/s influence, more performance metrics, as discussed in table 3, were developed to better represent quality and ranking of curves than TOI Importance AUC. To calculate AUC, we used the trapezoid function from the scipy python library.
A.5 Note on Model Selection Strategy: Minimal Supervision Framework
The CLARA pipeline employs unsupervised anomaly detection using URFs trained with synthetic light curves as the outlier class. Crucially, URF models are never exposed to labeled data with marked TOIs during training or anomaly scoring.
However, to interpret and steer model behavior, we introduce a light supervisory step using a small, representative benchmark - a randomly sampled subset of  4,000 real light curves from a sector (sector 2 in our case), enriched with known TOIs. This benchmark enables computation of an $\alpha$-score that summarizes each model’s alignment with chosen scientific priorities.
Once performance trends are established across URF subvariants using this small labeled set, models selected via $\alpha$ demonstrate robust and predictable behavior across larger datasets and different sectors - despite never being trained on those datasets. This approach combines the scalability of unsupervised learning with the interpretability of minimal supervision, enabling researchers to build reliable, scientifically aligned anomaly detectors without requiring extensive labeling efforts.
Appendix B Astrometric and Stellar Features Description
  • plx (Parallax): The apparent shift in a star’s position due to Earth’s orbital motion, measured in milliarcseconds (mas). Inversely proportional to distance ($d=1/\texttt{plx}$ in arcseconds).
  • pmra, pmdec (Proper Motion in RA/Dec): The motion of a star across the sky in right ascension and declination directions, respectively, measured in mas/yr. Reflects tangential velocity projected onto the sky.
  • Teff (Effective Temperature): The surface temperature of a star assuming it behaves like a blackbody, expressed in Kelvin. Influences stellar color and spectral type.
  • rad (Radius): The stellar radius, typically in solar radii ($R_{\odot}$). Derived from luminosity and temperature using Stefan–Boltzmann law.
  • mass (Mass): The estimated mass of the star, usually in solar masses ($M_{\odot}$), often inferred from stellar models.
  • lum (Luminosity): Total energy output of the star per unit time, expressed in solar luminosities ($L_{\odot}$). Strongly depends on both radius and temperature.
  • logg (Surface Gravity): Logarithm of the surface gravity ($g$) of the star, in cgs units. Sensitive to both mass and radius; higher for compact stars.
  • $v_{\text{tan}}$ (Tangential Velocity): The transverse velocity of the star relative to the Sun, calculated from proper motion and distance. Expressed in km/s.
  • ruwe (Renormalized Unit Weight Error): A goodness-of-fit statistic from Gaia astrometry; elevated values often indicate binarity, excess motion, or unresolved sources. Used as a quality indicator for Gaia solutions.
Appendix C Tables and Data
Table 8: Comparison of detection efficiency and enrichment of various transit detection pipelines. Detection rate is defined as the percentage of transit-like signals (confirmed or candidate) identified among the total light curves analyzed.
Method Light Curves per Sector Total Light Curves Project Candidates Detection Rate
CLARA Pipeline (till DTW-based DBSCAN clustering) $\sim 90$ 453 19 4.20%
DPMM Cluster 2 $\sim 23$ 114 16 14.04%
TESS SPOC (All Sectors, 2025) $\sim 20,000$ 1,680,000 (84 sectors) 7658 0.4569%
TESS-SPOC FFI Targets (Doyle et al., 2024) 2,890,583 7658 0.2650%
Appendix D Candidates from SPOC sector 1
Figure 35: Overlay of confirmed planets (in red on the right) from TESS SPOC Sector 1 in the high N-WRSS cluster
Table 9: Data on the 5 curves shown in appendix from planet-like class of sector 1
TIC ID Planet Name Period (days) Mass ($M_{\mathrm{J}}$) Radius ($R_{\mathrm{J}}$) Stellar Type Discovery Status
149603524 WASP-62b $4.411953\pm 3\times 10^{-6}$ $1.39\pm 0.06$ G-type Confirmed
144065872 WASP-95b $2.18466\pm 2\times 10^{-5}$ 1.44 $1.23\pm 0.06$ G-type Confirmed (2014)
260609205 TESS Candidate 4.4616249 2.83 Evolved (1.19 $M_{\odot}$) TESS TCE
290131778 HD 202772A b 3.3 $1.017^{+0.070}_{-0.068}$ $1.545^{+0.052}_{-0.060}$ Mildly evolved Confirmed
92352620 2.0 0.618 1.26 (est.) Confirmed (2014)
Table 10: SIMBAD label frequencies for the weakly supervised anomaly subset retrieved via MAST coordinate cross-matching. Labels are grouped into broader astrophysical categories for downstream classification.
SIMBAD Label Count Label Group
* 230 Stellar
Pl? 119 Planet-like
PM* 113 Stellar
err 79 Outlier
Pl 68 Planet-like
SB* 23 Binary Star
* 6 Binary Star
EB* 6 Binary Star
Er* 2 Stellar
s*b 2 Planet-like
LM* 1 Binary Star
BD* 1 Planet-like